### v2

# A quantitative Borg-Levinson theorem for a large class of unbounded potentials

Abstract.
We prove a quantitative Borg-Levinson theorem for a large class of unbounded potentials. We give a detailed proof when the dimension of the space is greater than or equal to five. We also indicate the modifications necessary to cover lower dimensions. In the last section, we briefly show how to extend our result to the anisotropic case.
Key words and phrases:
Borg-Levinson theorem, unbounded potentials, Dirichlet-to-Neumann map, resolvent estimate.
2010 Mathematics Subject Classification:
35R30
1. Introduction
Let $\Omega$ be a bounded $C^{1,1}$-domain of $\mathbb{R}^{n}$, $n\geq 5$, with boundary $\Gamma$. Let $p=2n/(n+2)$ and $q=2n/(n+4)$. Their respective conjugates are $p^{\prime}=2n/(n-2)$ and $q^{\prime}=2n/(n-4)$. That is we have $1/p+1/p^{\prime}=1$ and $1/q+1/q^{\prime}=1$.
Define on $H^{1}(\Omega)\times H^{1}(\Omega)$ the sesquilinear form $\mathfrak{a}_{V}$ associated with $V\in L^{n/2}(\Omega,\mathbb{R})$ by
$\mathfrak{a}_{V}(u,v)=\int_{\Omega}\left(\nabla u\cdot\nabla\overline{v}+Vu\overline{v}\right)dx.$
We prove (see Appendix A) that $\mathfrak{a}_{V}$ is bounded and coercive. Therefore, the bounded operator $A_{V}:H_{0}^{1}(\Omega)\rightarrow H^{-1}(\Omega)$ defined by
$\langle A_{V}u,v\rangle=\mathfrak{a}_{V}(u,v),\quad u,v\in H_{0}^{1}(\Omega)$
is self-adjoint (with respect to the pivot space $L^{2}(\Omega)$) and coercive. By applying [19, Theorem 2.37], we deduce that the spectrum of $A_{V}$, denoted $\sigma(A_{V})$, consists of a sequence $(\lambda_{V}^{k})$ satisfying
$\infty<\lambda_{V}^{1}\leq\lambda_{V}^{2}\leq\ldots\leq\lambda_{V}^{k}\leq\ldots$
and
$\lambda_{V}^{k}\rightarrow\infty\quad\mbox{as}\;k\rightarrow\infty.$
Moreover, $L^{2}(\Omega)$ admits an orthonormal basis $(\phi_{V}^{k})$ of eigenfunctions, each $\phi_{V}^{k}$ being associated with $\lambda_{V}^{k}$. That is, we have $\phi_{V}^{k}\in H_{0}^{1}(\Omega)$ and, for all $k\geq 1$ and $v\in H_{0}^{1}(\Omega)$,
$\int_{\Omega}\left(\nabla\phi_{V}^{k}\cdot\nabla\overline{v}+V\phi_{V}^{k}\overline{v}\right)dx=\lambda_{V}^{k}\int_{\Omega}\phi_{V}^{k}\overline{v}dx.$
In particular, $(-\Delta+V)\phi_{V}^{k}=\lambda_{V}^{k}\phi_{V}^{k}$ in the distributional sense, and since $(\lambda_{V}^{k}-V)\phi_{V}^{k}\in L^{q}(\Omega)$, we obtain from the usual $W^{2,q}$ regularity that $\phi_{V}^{k}\in W^{2,q}(\Omega)$.
Let $\gamma_{1}$ be the bounded operator from $W^{2,q}(\Omega)$ onto $W^{1-1/q,q}(\Gamma)$ given by
$\gamma_{1}w:=\partial_{\nu}w,\quad w\in C^{\infty}(\overline{\Omega}),$
where $\partial_{\nu}$ represents the derivative along the unitary exterior normal vector field $\nu$.
For simplicity, we hereafter use the notation
$\psi_{V}^{k}=\gamma_{1}\phi_{V}^{k},\quad k\geq 1,\;V\in L^{n/2}(\Omega,\mathbb{R}).$
The following Borg-Levinson type theorem was proved in [22].
Theorem 1.1.
Let $\ell\geq 1$ be an integer and $V_{1},V_{2}\in L^{m}(\Omega,\mathbb{R})$, where $m=n/2$ when $n\geq 4$ and $m>n/2$ when $n=3$, satisfy
$\lambda_{V_{1}}^{k}=\lambda_{V_{2}}^{k},\quad\psi_{V_{1}}^{k}=\psi_{V_{2}}^{k},\quad k\geq\ell.$
Then $V_{1}=V_{2}$.
Let $\Omega_{0}$ be an arbitrary neighborhood of $\Gamma$ in $\overline{\Omega}$ and $\Gamma_{0}$ be a nonempty open subset of $\Gamma$. Using unique continuation from Cauchy data for elliptic equations, we obtain the following consequence of theorem 1.1.
Corollary 1.1.
Let $\ell\geq 1$ be an integer and $V_{1},V_{2}\in L^{m}(\Omega,\mathbb{R})$, where $m=n/2$ when $n\geq 4$ and $m>n/2$ when $n=3$, satisfy $V_{1}{{}_{|\Omega_{0}}}\in L^{n}(\Omega_{0})$, $V_{1}=V_{2}$ in $\Omega_{0}$ and
$\lambda_{V_{1}}^{k}=\lambda_{V_{2}}^{k},\quad\psi_{V_{1}}^{k}{{}_{|\Gamma_{0}}}=\psi_{V_{2}}^{k}{{}_{|\Gamma_{0}}},\quad k\geq\ell.$
Then $V_{1}=V_{2}$.
We aim to establish a quantitative version of Theorem 1.1 when $\ell=1$. Before stating this result precisely, we need to introduce some definitions and notations. Fix $V_{0}\in L^{n/2}(\Omega,\mathbb{R})$ nonnegative and non identically equal to zero. Then let
$\mathscr{V}=\left\{V\in L^{n/2}(\Omega,\mathbb{R});\;|V|\leq V_{0}\right\}.$
Define for $V_{1},V_{2}\in\mathscr{V}$
$\mathscr{D}(V_{1},V_{2})=\sum_{k\geq 1}k^{-2/n}\left[|\lambda_{V_{1}}^{k}-\lambda_{V_{2}}^{k}|+\|\psi_{V_{2}}^{k}-\psi_{V_{1}}^{k}\|_{L^{q}(\Gamma)}\right],$
$\mathscr{D}_{+}(V_{1},V_{2})=\sum_{k\geq 1}k^{-2/n}\left[|\lambda_{V_{1}}^{k}-\lambda_{V_{2}}^{k}|+\|\phi_{V_{2}}^{k}-\phi_{V_{1}}^{k}\|_{W^{2,q}(\Omega)}\right].$
Note that, as the trace map $w\in W^{2,q}(\Omega)\mapsto\gamma_{1}w\in L^{q}(\Gamma)$ is bounded, if $\mathscr{D}_{+}(V_{1},V_{2})<\infty$, then $\mathscr{D}(V_{1},V_{2})<\infty$.
Next, fix $W_{0}\in L^{n}(\Omega,\mathbb{R})$ nonnegative and non identically equal to zero and set
$\mathscr{W}=\{(V_{1},V_{2})\in\mathscr{V}\times\mathscr{V};\;|V_{1}-V_{2}|\leq W_{0}\},$
$\mathscr{W}_{+}=\{(V_{1},V_{2})\in\mathscr{W};\;\mathscr{D}_{+}(V_{1},V_{2})<\infty\}.$
For fixed $t\in(1+1/q,2)$ we set
$\sigma:=\frac{12}{2-t},\quad\beta:=\frac{2}{(\sigma+n)(2n+4)+9(n+2)}.$
Theorem 1.2.
For all $(V_{1},V_{2})\in\mathscr{W}_{+}$ we have
(1.1) $\|V_{1}-V_{2}\|_{H^{-1}(\Omega)}\leq\mathbf{c}\mathscr{D}(V_{1},V_{2})^{\beta},$
where $\mathbf{c}=\mathbf{c}(n,\Omega,V_{0},W_{0},t)>0$ is a constant.
Remark 1.1.
Using Hölder’s inequality, we see that, in the definition of $\mathscr{W}_{+}$, $\mathscr{D}_{+}(V_{1},V_{2})<\infty$ can be replaced by
$\dot{\mathscr{D}}_{+}(V_{1},V_{2})=\sum_{k\geq 1}\left[|\lambda_{V_{1}}^{k}-\lambda_{V_{2}}^{k}|+\|\phi_{V_{2}}^{k}-\phi_{V_{1}}^{k}\|_{W^{2,q}(\Gamma)}\right]^{\eta}<\infty,$
for some $1\leq\eta<n/(n-2)$. Therefore, $\mathscr{D}(V_{1},V_{2})$ in (1.1) can be replaced by
$\dot{\mathscr{D}}(V_{1},V_{2})=\sum_{k\geq 1}\left[|\lambda_{V_{1}}^{k}-\lambda_{V_{2}}^{k}|+\|\psi_{V_{2}}^{k}-\psi_{V_{1}}^{k}\|_{L^{q}(\Gamma)}\right]^{\eta}.$
The proof of Theorem 1.2 is inspired by the previous work of the author and Stefanov [11] concerning the case of bounded potentials. However, we need in the present case more precise spectral estimates in $W^{2,q}(\Omega)$. Some parts of the Hilbertian analysis used in [11] are no longer valid for dealing with the case of unbounded potentials.
For the sake of clarity, we have intentionally excluded $n\in\{2,3,4\}$ because in this case some modifications are necessary. First, we need to replace $\mathscr{V}$ by
$\tilde{\mathscr{V}}=\{V\in L^{r}(\Omega,\mathbb{R});\;|V|\leq\tilde{V}_{0}\},$
where $r>\max(2,n/2)$ and $\tilde{V}_{0}\in L^{r}(\Omega)$ are arbitrarily fixed, with $\tilde{V}_{0}$ nonnegative and non identically equal to zero. With this modification, we must replace $q$ by $\tilde{q}=2r/(2+r)\in(1,2)$ (and then $\tilde{q}^{\prime}=2r/(r-2)$).
Theorem 1.2 still holds when $n\in\{3,4\}$ with $1+1/\tilde{q}<t<2$, provided we replace $\mathscr{D}(V_{1},V_{2})$, $\mathscr{D}_{+}(V_{1},V_{2})$, $\mathscr{W}$ and $\mathscr{W}_{+}$ respectively by
$\tilde{\mathscr{D}}(V_{1},V_{2})=\sum_{k\geq 1}k^{-2/n}\left[|\lambda_{V_{1}}^{k}-\lambda_{V_{2}}^{k}|+\|\psi_{V_{2}}^{k}-\psi_{V_{1}}^{k}\|_{L^{\tilde{q}}(\Gamma)}\right],$
$\tilde{\mathscr{D}}_{+}(V_{1},V_{2})=\sum_{k\geq 1}k^{-2/n}\left[|\lambda_{V_{1}}^{k}-\lambda_{V_{2}}^{k}|+\|\phi_{V_{2}}^{k}-\phi_{V_{1}}^{k}\|_{W^{2,q}(\Omega)}\right]$
$\tilde{\mathscr{W}}=\{(V_{1},V_{2})\in\tilde{\mathscr{V}}\times\tilde{\mathscr{V}};\;|V_{1}-V_{2}|\leq W_{0}\},$
$\tilde{\mathscr{W}}_{+}=\{(V_{1},V_{2})\in\tilde{\mathscr{W}};\;\tilde{\mathscr{D}}_{+}(V_{1},V_{2})<\infty\}.$
In the two-dimensional case, additional modifications are required. By using that $H^{1}(\Omega)$ is continuously embedded in $L^{2r/(r-1)}(\Omega)$, $p$ and $p^{\prime}$ must be replaced by $\tilde{p}=2r/(r+1)$ and $\tilde{p}^{\prime}=2r/(r-1)$, respectively. This choice ensures that Lemma 2.1 below remains valid in the two-dimensional case. On the other hand, the definition of $\mathscr{W}$ must be modified in order to ensure the validity of Lemma 2.7 with $q$ replaced by $\tilde{q}$. We then replace $\mathscr{W}$ by
$\hat{\mathscr{W}}=\{(V_{1},V_{2})\in\tilde{\mathscr{V}}\times\tilde{\mathscr{V}};\;|V_{1}-V_{2}|\leq\hat{W}_{0}\},$
where $\hat{W}_{0}\in L^{2r}(\Omega)$ is non-negative and not identically equal to zero, and we modify $\mathscr{W}_{+}$ accordingly. Under these modifications, Theorem 1.2 with $1+1/\tilde{q}<t<2$ is also valid in dimension two, provided that $\mathscr{D}(V_{1},V_{2})$, $\mathscr{D}_{+}(V_{1},V_{2})$, $\mathscr{W}$ and $\mathscr{W}_{+}$ are replaced by $\tilde{\mathscr{D}}(V_{1},V_{2})$, $\tilde{\mathscr{D}}_{+}(V_{1},V_{2})$, $\hat{\mathscr{W}}$ and $\hat{\mathscr{W}}_{+}$.
In the literature, the sequence $(\lambda_{V}^{k},\psi_{V}^{k})_{k\geq 1}$ (resp. $(\lambda_{V}^{k},\psi_{V}^{k})_{k\geq\ell}$) associated with $V\in L^{n/2}(\Omega,\mathbb{R})$ is called complete boundary spectral data (resp. incomplete boundary spectral data).
The first uniqueness result for potentials in $L^{r}$, $r>n/2$, is due to Pävärinta and Serov [21] in the case of complete boundary spectral data. This result was generalized by Pohjola [22] as shown by Theorem 1.1. The case of Robin’s boundary condition has recently been studied by the author, Metidji and Soccorsi [9] for potentials of the same class as in Theorem 1.1 in the case of incomplete boundary spectral data, precisely from $((\lambda_{V}^{k})_{k\geq\ell},(\psi_{V}^{k})_{k\geq 1})$. Bellassoued, Kian, Mannoubi and Soccorsi [6] established a result for potentials in $L^{\max(2,3n/5)}$ from the knowledge of asymptotic spectral data. For a precise definition of asymptotic spectral data, we refer to [16] where this notion was first introduced. A quantitative version of the uniqueness result in [6] can be found in [18]. A stability inequality, based on asymptotic spectral data in the case of a Robin’s boundary condition, was obtained by the author, Metidji, and Soccorsi in [10].
The first multidimensional Borg-Levinson type theorem was proved by Nachman, Sylvester and Uhlmann [20] by establishing the relationship between the complete boundary spectral data and the Dirichlet to Neumann map. The quantitative version of this result was proved by Alessandrini and Sylvester [1] by relating the original problem to a hyperbolic Dirichlet to Neumann map (see also [7] where the stability inequality is reformulated in an appropriate topology). The generalization of the result in [20] to the case of incomplete boundary spectral data is due to Isozaki [14, 15]. He used ideas borrowed from the Born approximation.
The magnetic case was first studied by Kian in [17] in the isotropic case and later generalized to the anisotropic case in [2]. For other results we refer, without being exhaustive, to [3, 4, 5, 8, 23, 24].
2. Preliminaries
This section presents the preliminary results that will be used in the following section to prove Theorem 1.2. First, since $H^{1}(\Omega)$ is continuously embedded in $L^{p^{\prime}}(\Omega)$, we define
$\mathbf{e}=\sup\left\{\|w\|_{L^{p^{\prime}}(\Omega)};\;u\in H^{1}(\Omega),\;\|u\|_{H^{1}(\Omega)}=1\right\}.$
The space $H_{0}^{1}(\Omega)$ will be endowed with the norm $\|\nabla\cdot\|_{L^{2}(\Omega)}$, and we set
$\mathbf{p}=\sup\left\{\|u\|_{H^{1}(\Omega)};\;u\in H_{0}^{1}(\Omega),\;\|\nabla u\|_{L^{2}(\Omega)}=1\right\}.$
2.1. Weyl’s asymptotic formula
We have the following improvement of [22, Lemma 5.2].
Lemma 2.1.
Let $V\in\mathscr{V}$ and $u\in H_{0}^{1}(\Omega)$. Then for any $\epsilon>0$, there exists a constant $\mathbf{c}_{\epsilon}=\mathbf{c}_{\epsilon}(n,\Omega,V_{0},\epsilon)>0$ such that
(2.1) $\|Vu^{2}\|_{L^{1}(\Omega)}\leq\epsilon\|\nabla u\|_{L^{2}(\Omega)}+\mathbf{c}_{\epsilon}\|u\|_{L^{2}(\Omega)}.$
Proof.
By applying Hölder’s inequality, we obtain
$\|Vu^{2}\|_{L^{1}(\Omega)}\leq\|V_{0}u^{2}\|_{L^{1}(\Omega)}$ $\leq\|V_{0}u\|_{L^{p}(\Omega)}\|u\|_{L^{p^{\prime}}(\Omega)}$
$\leq\mathbf{p}\mathbf{e}\|V_{0}u\|_{L^{p}(\Omega)}\|\nabla u\|_{L^{2}(\Omega)}.$
Then, we have for all $\epsilon>0$
(2.2) $\|Vu^{2}\|_{L^{1}(\Omega)}\leq 2\mathbf{p}\mathbf{e}\epsilon\|\nabla u\|_{L^{2}(\Omega)}^{2}+2\mathbf{p}\mathbf{e}\epsilon^{-1}\|V_{0}u\|_{L^{p}(\Omega)}^{2}.$
On the other hand, there exists a constant $\mathbf{c_{0}}=\mathbf{c}_{0}(n,\Omega)>0$ so that for all $W_{0}\in C_{0}^{\infty}(\Omega)$
$\|V_{0}u\|_{L^{p}(\Omega)}^{2}$ $\leq 2\|W_{0}u\|_{L^{p}(\Omega)}^{2}+2\|(V_{0}-W_{0})u\|_{L^{p}(\Omega)}^{2}$
$\leq 2\|W_{0}\|_{L^{\infty}(\Omega)}^{2}\|u\|_{L^{p}(\Omega)}^{2}+2\|(V_{0}-W_{0})\|_{L^{n/2}(\Omega)}^{2}\|u\|_{L^{p^{\prime}}(\Omega)}^{2}$
$\leq\mathbf{c_{0}}\|W_{0}\|_{L^{\infty}(\Omega)}^{2}\|u\|_{L^{2}(\Omega)}^{2}+\mathbf{c_{0}}\|(V_{0}-W_{0})\|_{L^{n/2}(\Omega)}^{2}\|u\|_{L^{p^{\prime}}(\Omega)}^{2}$
$\leq\mathbf{c_{0}}\|W_{0}\|_{L^{\infty}(\Omega)}^{2}\|u\|_{L^{2}(\Omega)}^{2}+\mathbf{c_{0}}\|(V_{0}-W_{0})\|_{L^{n/2}(\Omega)}^{2}\|\nabla u\|_{L^{2}(\Omega)}^{2}.$
As $C_{0}^{\infty}(\Omega)$ is dense in $L^{n/2}(\Omega)$, we choose $W_{0}$ in such a way to obtain
(2.3) $\|V_{0}u\|_{L^{p}(\Omega)}^{2}\leq\mathbf{c_{0}}\|W_{0}\|_{L^{\infty}(\Omega)}^{2}\|u\|_{L^{2}(\Omega)}^{2}+\mathbf{c_{0}}\epsilon^{2}\|\nabla u\|_{L^{2}(\Omega)}^{2}.$
This and (2.2) imply
(2.4) $\|Vu^{2}\|_{L^{1}(\Omega)}\leq 2(\mathbf{c}_{0}+1)\mathbf{p}\mathbf{e}\epsilon\|\nabla u\|_{L^{2}(\Omega)}^{2}+2\mathbf{c}_{0}\mathbf{p}\mathbf{e}\epsilon^{-1}\|W_{0}\|_{L^{\infty}(\Omega)}^{2}\|u\|_{L^{2}(\Omega)}^{2},$
from which we deduce the expected inequality. ∎
Proposition 2.1.
There exist two constants $\mathfrak{c}_{0}=\mathfrak{c}_{0}(n,\Omega)\geq 1$ and $\mathfrak{c}_{1}=\mathfrak{c}_{1}(n,\Omega,V_{0})>0$ such that for all $V\in\mathscr{V}$ we have
(2.5) $\mathfrak{c}_{0}^{-1}k^{2/n}-\mathfrak{c}_{1}\leq\lambda_{V}^{k}\leq\mathfrak{c}_{0}k^{2/n}+\mathfrak{c}_{1},\quad k\geq 1.$
Proof.
In this proof, $\mathfrak{c}_{1}=\mathfrak{c}_{1}(n,\Omega,V_{0})>0$ is a generic constant. Let $V\in\mathscr{V}$. By applying (2.1) with $\epsilon=1$, we find
(2.6) $\mathfrak{a}_{V}(u,u)\leq 2\|\nabla u\|_{L^{2}(\Omega)}+\mathfrak{c}_{1}\|u\|_{L^{2}(\Omega)},\quad u\in H_{0}^{1}(\Omega).$
On the other hand, we have
$\mathfrak{a}_{V}(u,u)\geq\mathfrak{a}_{-V_{0}}(u,u)=\|\nabla u\|_{L^{2}(\Omega)}-\|V_{0}u^{2}\|_{L^{1}(\Omega)},\quad u\in H_{0}^{1}(\Omega).$
By applying again (2.1) with $\epsilon=1/2$, we get
$\mathfrak{a}_{V}(u,u)\geq\frac{1}{2}\|\nabla u\|_{L^{2}(\Omega)}-\mathfrak{c}_{1}\|u\|_{L^{2}(\Omega)},\quad u\in H_{0}^{1}(\Omega).$
Let $(\mu_{k})$ denotes the non decreasing sequence of eigenvalues of $\mathfrak{a}_{0}$ (that is $\mathfrak{a}_{V}$ with $V=0$). Then the min-max principle yields
$\frac{\mu_{k}}{2}-\mathfrak{c}_{1}\leq\lambda_{V}^{k}\leq 2\mu_{k}+\mathfrak{c}_{1},\quad k\geq 1.$
We complete the proof using Weyl’s classical asymptotic formula saying that $\mu_{k}\sim k^{2/n}$ as $k$ tends to $\infty$. ∎
2.2. Dirichlet to Neumann map
Let $\rho(A_{V})=\mathbb{C}\setminus\sigma(A_{V})$, usually called the resolvent set of $A_{V}$. We consider the BVP
(2.7) $\left\{\begin{array}[]{ll}(-\Delta+V-\lambda)u=0\quad&\mbox{in}\;\Omega,\\ u=\varphi&\mbox{on}\;\Gamma.\end{array}\right$.
Let $V\in L^{n/2}(\Omega,\mathbb{R})$ and $\lambda\in\rho(A_{V})$. It follows from Theorem B.1 in Appendix B that, for all $\varphi\in W^{2-1/q,q}(\Gamma)$, the BVP (2.7) admits a unique solution $u_{V}(\lambda)(\varphi)\in W^{2,q}(\Omega)$. Furthermore, we have
(2.8) $\|u_{V}(\lambda)(\varphi)\|_{W^{2,q}(\Omega)}\leq\varkappa\|\varphi\|_{W^{2-1/q,q}(\Gamma)},$
where $\varkappa=\varkappa(n,\Omega,V,\lambda)>0$ is a constant.
Define the family of Dirichlet to Neumann maps $(\Lambda_{V}(\lambda))_{\lambda\in\rho(A_{V})}$ associated with $V\in L^{n/2}(\Omega,\mathbb{R})$ as follows
$\Lambda_{V}(\lambda):\varphi\in W^{2-1/q,q}(\Gamma)\mapsto\gamma_{1}u_{V}(\lambda)(\varphi)\in W^{1-1/q,q}(\Gamma).$
By the continuity of the trace map $w\in W^{2,q}(\Omega)\mapsto\gamma_{1}w\in W^{1-1/q,q}(\Gamma)$, it follows from (2.8) that the mapping $\Lambda_{V}(\lambda)$ defines a bounded operator.
2.3. Analyticity of solutions
In this subsection, $\mathbf{c}=\mathbf{c}(n,\Omega,V_{0})>0$ denotes a generic constant. Let $V\in\mathscr{V}$ and $k\geq 1$. Since
$\|\nabla\phi_{V}^{k}\|_{L^{2}(\Omega)}^{2}\leq|\lambda_{V}^{k}|+\|V|\phi_{V}^{k}|^{2}\|_{L^{1}(\Omega)},$
(2.1) with $\epsilon=1/2$ yields
$\|\nabla\phi_{V}^{k}\|_{L^{2}(\Omega)}\leq\mathbf{c}(1+|\lambda_{V}^{k}|)^{1/2}$
and then
(2.9) $\|\phi_{V}^{k}\|_{L^{p^{\prime}}(\Omega)}\leq\mathbf{c}(1+|\lambda_{V}^{k}|)^{1/2}.$
In consequence, we have
(2.10) $\|V\phi_{V}^{k}\|_{L^{p}(\Omega)}\leq\mathbf{c}(1+|\lambda_{V}^{k}|)^{1/2}.$
By using that $\Delta\phi_{V}^{k}=(\lambda_{V}^{k}-V)\phi_{V}^{k}\in L^{q}(\Omega)$, we get $\phi_{k}\in W^{2,q}(\Omega)$. In light of $W^{2,q}$ a priori estimate (e.g. [13, Theorem 2.3.3.6]), (2.10) yields
(2.11) $\|\phi_{k}\|_{W^{2,q}(\Omega)}\leq\mathbf{c}(1+|\lambda_{V}^{k}|).$
Let $\lambda\in\rho(A_{V})$, $\varphi\in W^{2-1/q,q}(\Gamma)$ and $u:=u_{V}(\lambda)(\varphi)$. Then
$0=\int_{\Omega}(-\Delta u+(V-\lambda)u)\overline{\phi_{V}^{k}}dx.$
Let $(\cdot,\cdot)$ denotes the inner product of $L^{2}(\Omega)$. By performing an integration by parts, we derive from the preceding identity
(2.12) $(u,\phi_{V}^{k})=-\frac{\langle\varphi,\psi_{V}^{k}\rangle}{\lambda_{V}^{k}-\lambda}.$
Here and henceforth,
$\langle\varphi,\psi_{V}^{k}\rangle:=\int_{\Gamma}\varphi\overline{\psi_{V}^{k}}d\sigma.$
Inequality (2.12) then yields
$u_{V}(\lambda)(\varphi)=-\sum_{k\geq 1}\frac{\langle\varphi,\psi_{V}^{k}\rangle}{\lambda_{V}^{k}-\lambda}\phi_{k}.$
In particular, we have
(2.13) $\sum_{k\geq 1}\left|\frac{\langle\varphi,\psi_{V}^{k}\rangle}{\lambda_{V}^{k}-\lambda}\right|^{2}=\|u_{V}(\lambda)(g)\|_{L^{2}(\Omega)}^{2}.$
Let $\lambda_{0}\in\rho(A_{V})$ and $d=\mathrm{dist}(\lambda_{0},\sigma(A_{V}))$. Since
$|\lambda_{k}-\lambda|\geq|\lambda_{k}-\lambda_{0}|-d/2,\quad\lambda\in B(\lambda_{0},d/2),$
and $|\lambda_{k}-\lambda_{0}|-d/2\sim|\lambda_{k}-\lambda_{0}|$ when $k$ goes to $\infty$, by using (2.13) we derive that the series $\sum_{k\geq 1}\frac{\langle\varphi,\psi_{V}^{k}\rangle}{\lambda_{V}^{k}-\lambda}\phi_{k}$ converges in $L^{2}(\Omega)$ uniformly in $B(\lambda_{0},d/2)$ and therefore in any compact subset of $\rho(A_{V})$. Weierstrass’s theorem then allows us to conclude that the mapping
$\lambda\in\rho(A_{V})\mapsto u_{V}(\lambda)(\varphi)\in L^{2}(\Omega)$
is holomorphic and
(2.14) $u_{V}^{(j)}(\lambda)(\varphi)=-j!\sum_{k\geq 1}\frac{\langle\varphi,\psi_{V}^{k}\rangle}{(\lambda_{V}^{k}-\lambda)^{j+1}}\phi_{V}^{k},\quad j\geq 0.$
The following notation will be used hereinafter.
$w^{(j)}:=\frac{d^{j}}{d\lambda^{j}}w.$
Let $\mathscr{S}$ be the bounded affine function $f\in L^{q}(\Omega)\mapsto\mathscr{S}f:=w\in W^{2,q}(\Omega)$, where $w$ is the unique solution of the BVP
$\Delta w=f\;\mathrm{in}\;\Omega,\quad u=\varphi\;\mathrm{on}\;\Gamma.$
By observing that $u_{V}(\lambda)(\varphi)=\mathscr{S}((\lambda-V)u_{V}(\lambda)(\varphi))$, we conclude that $\lambda\in\rho(A_{V})\mapsto u_{V}(\lambda)(\varphi)\in W^{2,q}(\Omega)$ is holomorphic and by induction in $j\geq 1$, we have, where $u(\lambda):=u_{V}(\lambda)(\varphi)$,
(2.15) $(-\Delta+V-\lambda)u^{(j)}(\lambda)=ju^{(j-1)}(\lambda)\;\mathrm{in}\;\Omega,\quad u^{(j)}(\lambda)=0\;\mathrm{on}\;\Gamma.$
2.4. Resolvent estimates
Recall that the resolvent associated with the operator $A_{V}$ is defined as follows
$R_{V}(\lambda)=(A_{V}-\lambda)^{-1}:H^{-1}(\Omega)\rightarrow H_{0}^{1}(\Omega),\quad\lambda\in\rho(A_{V}).$
For $\in\mathbb{C}, z\Re z$ and $\Im z$ will respectively denote the real and imaginary parts of $z$. Let
$\Sigma_{0}=\{\lambda\in\mathbb{C};\;|\Re\lambda|\geq 1,\;|\Im\lambda|\geq 1,\;|\Re\lambda||\Im\lambda|^{-1}\geq 1/2\}.$
In all of this subsection, $\mathbf{c}=\mathbf{c}(n,\Omega,V_{0})>0$ will denote a generic constant.
Lemma 2.2.
Let $V\in\mathscr{V}$, $\lambda\in\Sigma_{0}$ and $f\in L^{2}(\Omega)$. Then
(2.16) $\|R_{V}(\lambda)f\|_{L^{2}(\Omega)}\leq|\Im\lambda|^{-1}\|f\|_{L^{2}(\Omega)},$
(2.17) $\|R_{V}(\lambda)f\|_{H_{0}^{1}(\Omega)}\leq\mathbf{c}|\Re\lambda|^{1/2}|\Im\lambda|^{-1}\|f\|_{L^{2}(\Omega)},$
(2.18) $\|R_{V}(\lambda)f\|_{H_{0}^{1}(\Omega)}\leq\mathbf{c}|\Re\lambda||\Im\lambda|^{-1}\|f\|_{L^{p}(\Omega)}.$
Proof.
Let $u:=R_{V}(\lambda)f$. Green’s formula then yields
$\int_{\Omega}|\nabla u|^{2}dx-\int_{\Omega}(\lambda-V)|u|^{2}dx=\int_{\Omega}f\overline{u}dx.$
Hence
(2.19) $\int_{\Omega}|\nabla u|^{2}dx-\int_{\Omega}(\Re\lambda-V)|u|^{2}dx=\Re\int_{\Omega}f\overline{u}dx,$
(2.20) $-\Im\lambda\int_{\Omega}|u|^{2}dx=\Im\int_{\Omega}f\overline{u}dx.$
Then (2.16) follows easily from (2.20).
Next, by using (2.19) and (2.1) with $\epsilon=1/2$, we get
$\|\nabla u\|_{L^{2}(\Omega)}^{2}\leq\mathbf{c}|\Re\lambda|\|u\|_{L^{2}(\Omega)}^{2}+\|f\overline{u}\|_{L^{1}(\Omega)}.$
This inequality combined with (2.20) yields
(2.21) $\|\nabla u\|_{L^{2}(\Omega)}^{2}$ $\leq\mathbf{c}|\Re\lambda||\Im\lambda|^{-1}\|f\overline{u}\|_{L^{1}(\Omega)}$
$\leq\mathbf{c}|\Re\lambda||\Im\lambda|^{-1}\|u\|_{L^{2}(\Omega)}\|f\|_{L^{2}(\Omega)}.$
By combining this inequality and (2.16), we obtain
$\|\nabla u\|_{L^{2}(\Omega)}^{2}\leq\mathbf{c}|\Re\lambda||\Im\lambda|^{-2}\|f\|_{L^{2}(\Omega)}^{2}.$
That is we proved (2.17).
On the other hand, applying Hölder’s inequality, we get
$\|f\overline{u}\|_{L^{1}(\Omega)}\leq\|u\|_{L^{p^{\prime}}(\Omega)}\|f\|_{L^{p}(\Omega)},$
and hence, for all $\epsilon>0$,
$\|f\overline{u}\|_{L^{1}(\Omega)}\leq\frac{\epsilon}{2}\|u\|_{L^{p^{\prime}}(\Omega)}^{2}+\frac{1}{2\epsilon}\|f\|_{L^{p}(\Omega)}^{2}.$
Therefore
(2.22) $\|f\overline{u}\|_{L^{1}(\Omega)}\leq\frac{\epsilon\mathbf{e}\mathbf{p}}{2}\|\nabla u\|_{L^{2}(\Omega)}^{2}+\frac{1}{2\epsilon}\|f\|_{L^{p}(\Omega)}^{2}.$
By using (2.22) in the first inequality of (2.21), we find
$\|\nabla u\|_{L^{2}(\Omega)}^{2}\leq c_{0}|\Re\lambda||\Im\lambda|^{-1}\left[\frac{\epsilon\mathbf{e}\mathbf{p}}{2}\|\nabla u\|_{L^{2}(\Omega)}^{2}+\frac{1}{2\epsilon}\|f\|_{L^{p}(\Omega)}^{2}\right],$
where $c_{0}>0$ is a universal constant. By taking in this inequality $\epsilon$ so that $c_{0}|\Re\lambda||\Im\lambda|^{-1}\epsilon\mathbf{e}\mathbf{p}=1$, we obtain (2.18). ∎
Define
$\Sigma=\{\lambda_{\tau}=(\tau+i)^{2};\;\tau\geq 2\}\;(\subset\Sigma_{0}).$
Corollary 2.1.
Let $V\in\mathscr{V}$, $\lambda_{\tau}=(\tau+i)^{2}\in\Sigma$ and $f\in L^{2}(\Omega)$. Then
(2.23) $2\|R_{V}(\lambda_{\tau})f\|_{L^{2}(\Omega)}\leq\tau^{-1}\|f\|_{L^{2}(\Omega)},$
(2.24) $\|R_{V}(\lambda_{\tau})f\|_{L^{p^{\prime}}(\Omega)}\leq\mathbf{c}\tau\|f\|_{L^{p}(\Omega)}.$
Define
(2.25) $p_{\theta}=2n/(n+2\theta),\quad p_{\theta}^{\prime}=2n/(n-2\theta),\quad\theta\in[0,1],$
and note that $p_{\theta}^{\prime}$ is the conjugate of $p_{\theta}$. By applying Riesz-Thorin’s theorem, we obtain from (2.23) and (2.24) the following result.
Proposition 2.2.
Let $V\in\mathscr{V}$. Then for all $\lambda_{\tau}=(\tau+i)^{2}\in\Sigma$, $R_{V}(\lambda_{\tau})$ maps continuously $L^{p_{\theta}}(\Omega)$ into $L^{p_{\theta}^{\prime}}(\Omega)$ and the following estimate holds
(2.26) $\|R_{V}(\lambda_{\tau})f\|_{L^{p_{\theta}^{\prime}}(\Omega)}\leq\mathbf{c}\tau^{-1+2\theta}\|f\|_{L^{p_{\theta}}(\Omega)},\quad f\in L^{p_{\theta}}(\Omega).$
Lemma 2.3.
Let $V\in\mathscr{V}$. There exists $\lambda_{\ast}=\lambda_{\ast}(n,\Omega,V_{0})\geq 2$ such that for any $\lambda\in\mathbb{C}\setminus\mathbb{R}$ satisfying $\Re\lambda\geq\lambda_{\ast}$ and $f\in L^{2}(\Omega)$ we have
(2.27) $\|R_{V}(\lambda)f\|_{L^{2}(\Omega)}\leq\sqrt{2}|\Re\lambda|^{-1/2}\|f\|_{L^{2}(\Omega)},$
(2.28) $\|R_{V}(\lambda)f\|_{H_{0}^{1}(\Omega)}\leq\sqrt{2}\|f\|_{L^{2}(\Omega)}.$
Proof.
Let $\lambda\in\mathbb{C}\setminus\mathbb{R}$ with $\Re\lambda\leq 0$, $f\in L^{2}(\Omega)$ and $u:=R_{V}(\lambda)f$. From (2.19), we obtain
(2.29) $\int_{\Omega}|\nabla u|^{2}dx+|\Re\lambda|\int_{\Omega}|u|^{2}dx=-\int_{\Omega}V|u|^{2}dx+\Re\int_{\Omega}f\overline{u}dx.$
From (2.1), we get
(2.30) $\|Vu^{2}\|_{L^{1}(\Omega)}\leq\frac{1}{2}\|\nabla u\|_{L^{2}(\Omega)}^{2}+\mathbf{c}\|u\|_{L^{2}(\Omega)}^{2}.$
On the other hand, we have
(2.31) $\Re\int_{\Omega}f\overline{u}dx\leq\|u\|_{L^{2}(\Omega)}^{2}+\|f\|_{L^{2}(\Omega)}^{2}.$
By combining (2.29), (2.30) and (2.31), we obtain
(2.32) $\int_{\Omega}|\nabla u|^{2}dx+2(|\Re\lambda|-1-\mathbf{c})\int_{\Omega}|u|^{2}dx\leq 2\|f\|_{L^{2}(\Omega)}^{2}.$
Let $\lambda_{\ast}=2(1+\mathbf{c})$. If $|\Re\lambda|\geq\lambda_{\ast}$, then (2.32) yields
$\|u\|_{L^{2}(\Omega)}\leq\sqrt{2}|\Re\lambda|^{-1/2}\|f\|_{L^{2}(\Omega)},$
$\|\nabla u\|_{L^{2}(\Omega)}\leq\sqrt{2}\|f\|_{L^{2}(\Omega)}.$
These are the expected inequalities. ∎
Recall that the trace operator $\gamma_{0}:W^{2,q}(\Omega)\rightarrow W^{2-1/q,q}(\Gamma)$ defined by $\gamma_{0}w=w_{|\Gamma}$, $w\in C^{\infty}(\overline{\Omega})$, is bounded. From [13, Theorem 2.3.3.6], there exists $c_{0}=c_{0}(n,\Omega)>0$ and $\mu_{0}=\mu_{0}(n,\Omega)>0$ such that for any $\mu\geq\mu_{0}$ and $w\in W^{2,q}(\Omega)$ we have
(2.33) $\|w\|_{W^{2,q}(\Omega)}\leq c_{0}\left(\|(\Delta-\mu)w\|_{L^{q}(\Omega)}+\|\gamma_{0}w\|_{W^{2-1/q,q}(\Gamma)}\right).$
Define
$\Pi_{1}=\{\lambda\in\mathbb{C}\setminus\mathbb{R};\;-\Re\lambda\geq\max(\mu_{0},\lambda_{\ast}),\;|\Im\lambda||\Re\lambda|^{-1/2}\leq 1\},$
where $\mu_{0}$ is as above and $\lambda^{\ast}$ is as in Lemma 2.3.
Lemma 2.4.
Let $V\in\mathscr{V}$, $f\in L^{2}(\Omega)$ and $\lambda\in\Pi_{1}$. Then
(2.34) $\|R_{V}(\lambda)(f)\|_{W^{2,q}(\Omega)}\leq\mathbf{c}\|f\|_{L^{2}(\Omega)}.$
Proof.
Let $u:=R_{V}(\lambda)f\in W^{2,q}(\Omega)$, $v:=\Re u$ and $w:=\Im u$. As $\Delta u=(\lambda-V)u+f$, we obtain
$(\Delta+\Re\lambda)v=\Im\lambda w+Vv-\Re f,\quad(\Delta+\Re\lambda)w=-\Im\lambda v+Vw-\Im f.$
Then (2.33) yields
$\|v\|_{W^{2,q}(\Omega)}$ $\leq c_{0}\|(\Delta+\Re\lambda)v\|_{L^{q}(\Omega)}$
$\leq c_{0}\|\Im\lambda w+Vv-\Re f\|_{L^{q}(\Omega)}.$
As
$\|Vv\|_{L^{q}(\Omega)}\leq\|V_{0}\|_{L^{n/2}(\Omega)}\|v\|_{L^{2}(\Omega)},$
we obtain
$\|v\|_{W^{2,q}(\Omega)}\leq\mathbf{c}\left(|\Im\lambda|\|w\|_{L^{2}(\Omega)}+\|v\|_{L^{2}(\Omega)}+\|f\|_{L^{2}(\Omega)}\right).$
Similarly, we have
$\|w\|_{W^{2,q}(\Omega)}\leq\mathbf{c}\left(|\Im\lambda|\|v\|_{L^{2}(\Omega)}+\|w\|_{L^{2}(\Omega)}+\|f\|_{L^{2}(\Omega)}\right).$
By combining the last two inequalities, we find
$\|u\|_{W^{2,q}(\Omega)}\leq\mathbf{c}\left(|\Im\lambda|\|u\|_{L^{2}(\Omega)}+\|u\|_{L^{2}(\Omega)}+\|f\|_{L^{2}(\Omega)}\right).$
This and (2.27) imply
$\|u\|_{W^{2,q}(\Omega)}\leq\mathbf{c}\left(|\Im\lambda||\Re\lambda|^{-1/2}+1\right)\|f\|_{L^{2}(\Omega)},$
from which (2.34) follows. ∎
2.5. The difference between two Dirichlet to Neumann maps
Let
$\mathcal{B}=W^{2-1/q,q}(\Gamma)\cap L^{q^{\prime}}(\Gamma)$
be endowed with the norm
$\|\varphi\|:=\|\varphi\|_{W^{2-1/q,q}(\Gamma)}+\|\varphi\|_{L^{q^{\prime}}(\Gamma)},\quad\varphi\in\mathcal{B}.$
As in the previous subsection, $\mathbf{c}=\mathbf{c}(n,\Omega,V_{0})>0$ still denotes a generic constant.
Let $V\in\mathscr{V}$, $\varphi\in\mathcal{B}$ and $\lambda\in\Pi_{1}$. We proceed as in the proof of Lemma 2.4 to obtain
(2.35) $\|u_{V}(\lambda)(\varphi)\|_{W^{2,q}(\Omega)}\leq\mathbf{c}\left(|\Im\lambda|\|u_{V}(\lambda)(\varphi)\|_{L^{2}(\Omega)}+\|\varphi\|\right).$
Lemma 2.5.
Let $V\in\mathscr{V}$ and $\varphi\in\mathcal{B}$. There exists $\mu_{\ast}=\mu_{\ast}(n,\Omega,V_{0})\geq\mu_{0}$ so that for all $\lambda\in\mathbb{C}\setminus\mathbb{R}$ with $\Re\lambda\geq\mu_{\ast}$ we have
(2.36) $\|u_{V}(\lambda)(\varphi)\|_{L^{2}(\Omega)}\leq\mathbf{c}|\Re\lambda|^{-1/2}\|u_{V}(\lambda)(\varphi)\|_{W^{2,q}(\Omega)}^{1/2}\|\varphi\|^{1/2},$
(2.37) $\|u_{V}(\lambda)(\varphi)\|_{L^{p^{\prime}}(\Omega)}\leq\mathbf{c}\|u_{V}(\lambda)(\varphi)\|_{W^{2,q}(\Omega)}^{1/2}\|\varphi\|^{1/2}.$
Proof.
Let $u:=u_{V}(\lambda)(\varphi)$. By applying Green’s formula, we find
$\int_{\Omega}|\nabla u|^{2}dx+\int_{\Omega}(V-\lambda)|u|^{2}dx=\int_{\Gamma}\gamma_{1}u\overline{\varphi}d\sigma.$
By proceeding as in the proof of Lemma 2.3, we find $\mu_{\ast}=\mu_{\ast}(n,\Omega,V_{0})\geq\mu_{0}$ such that for all $\lambda\in\mathbb{C}\setminus\mathbb{R}$ satisfying $\Re\lambda\geq\mu_{\ast}$ we have
$\|\nabla u\|_{L^{2}(\Omega)}^{2}+|\Re\lambda|\|u\|_{L^{2}(\Omega)}^{2}\leq\mathbf{c}\|\varphi\gamma_{1}u\|_{L^{1}(\Gamma)}$
and then
$\|\nabla u\|_{L^{2}(\Omega)}^{2}+|\Re\lambda|\|u\|_{L^{2}(\Omega)}^{2}$ $\leq\mathbf{c}\|\gamma_{1}u\|_{L^{q}(\Gamma)}\|\varphi\|_{L^{q^{\prime}}(\Gamma)}$
$\leq\mathbf{c}\|\gamma_{1}u\|_{L^{q}(\Gamma)}\|\varphi\|.$
The proof is completed by using that $\gamma_{1}$ is bounded from $W^{2,q}(\Omega)$ into $L^{q}(\Gamma)$. ∎
Define
$\Pi_{2}=\{\lambda\in\mathbb{C}\setminus\mathbb{R};\;-\Re\lambda\geq\max(\lambda_{\ast},\mu_{\ast}),\;|\Im\lambda||\Re\lambda|^{-1/2}\leq 1\}\;(\subset\Pi_{1}).$
where $\lambda_{\ast}$ is as in Lemma 2.3 and $\mu_{\ast}$ is as in Lemma 2.5.
Let $V\in\mathscr{V}$, $\lambda\in\Pi_{2}$ and $\varphi\in\mathcal{B}$. By combining (2.35) and (2.36), we derive
$\|u_{V}(\lambda)(\varphi)\|_{W^{2,q}(\Omega)}\leq\mathbf{c}\left(\|u_{V}(\lambda)(\varphi)\|_{W^{2,q}(\Omega)}^{1/2}\|\varphi\|^{1/2}+\|\varphi\|\right).$
Whence
$\|u_{V}(\lambda)(\varphi)\|_{W^{2,q}(\Omega)}\leq\frac{1}{2}\|u_{V}(\lambda)(\varphi)\|_{W^{2,q}(\Omega)}+\mathbf{c}\|\varphi\|$
and hence
(2.38) $\|u_{V}(\varphi)\|_{W^{2,q}(\Omega)}\leq\mathbf{c}\|\varphi\|.$
This and (2.36) imply
(2.39) $\|u_{V}(\lambda)(\varphi)\|_{L^{2}(\Omega)}\leq\mathbf{c}|\Re\lambda|^{-1/2}\|\varphi\|.$
Similarly, we obtain from (2.37)
(2.40) $\|u_{V}(\lambda)(\varphi)\|_{L^{p^{\prime}}(\Omega)}\leq\mathbf{c}\|\varphi\|.$
For all non negative integer $j$, $\mathbf{c}_{j}=\mathbf{c}_{j}(n,\Omega,V_{0},j)>0$ will denote in the remaining part of this subsection a generic constant.
Lemma 2.6.
Let $j\geq 0$ be an integer, $V\in\mathscr{V}$, $\lambda\in\Pi_{2}$ and $\varphi\in\mathcal{B}$. Then we have
(2.41) $\|u_{V}^{(j)}(\varphi)(\lambda)(\varphi)\|_{W^{2,q}(\Omega)}\leq\mathbf{c}_{j}|\Re\lambda|^{-j/2}\|\varphi\|,$
(2.42) $\|u_{V}^{(j)}(\varphi)(\lambda)(\varphi)\|_{L^{2}(\Omega)}\leq\mathbf{c}_{j}|\Re\lambda|^{-(j+1)/2}\|\varphi\|,$
(2.43) $\|u_{V}^{(j)}(\varphi)(\lambda)(\varphi)\|_{L^{p^{\prime}}(\Omega)}\leq\mathbf{c}_{j}|\Re\lambda|^{-j/2}\|\varphi\|.$
Proof.
Let $u(\lambda):=u_{V}(\lambda)(\varphi)$. We have from (2.15)
$(-\Delta+V-\lambda)u^{(1)}(\lambda)=u(\lambda),\quad\gamma_{0}u^{(1)}=0.$
Thus,
$u^{(1)}(\lambda)=R_{V}(\lambda)(u(\lambda)).$
Whence, according to (2.34),
(2.44) $\|u^{(1)}(\lambda)\|_{W^{2,q}(\Omega)}\leq\mathbf{c}\|u(\lambda)\|_{L^{2}(\Omega)}.$
This and (2.39) yield
(2.45) $\|u^{(1)}(\lambda)\|_{W^{2,q}(\Omega)}\leq\mathbf{c}|\Re\lambda|^{-1/2}\|\varphi\|.$
On the other hand, it follows from (2.27) and (2.28)
$\|u^{(1)}(\lambda)\|_{L^{2}(\Omega)}\leq\sqrt{2}|\Re\lambda|^{-1/2}\|u(\lambda)\|_{L^{2}(\Omega)},$
$\|u^{(1)}(\lambda)\|_{L^{p^{\prime}}(\Omega)}\leq\sqrt{2}\|u(\lambda)\|_{L^{2}(\Omega)}.$
By applying (2.39) and (2.40), we find
(2.46) $\|u^{(1)}(\lambda)\|_{L^{2}(\Omega)}\leq\mathbf{c}|\Re\lambda|^{-2(1/2)}\|\varphi\|,$
(2.47) $\|u^{(1)}(\lambda)\|_{L^{p^{\prime}}(\Omega)}\leq\mathbf{c}|\Re\lambda|^{-1/2}\|\varphi\|.$
We get again from (2.15)
$(-\Delta+V-\lambda)u^{(2)}(\lambda)=2u^{(1)}(\lambda),\quad\gamma_{0}u^{(2)}=0.$
We repeat the previous calculations replacing $u(\lambda)$ with $u^{(1)}(\lambda)$ in order to obtain
$\|u^{(2)}(\lambda)\|_{W^{2,q}(\Omega)}\leq\mathbf{c}|\Re\lambda|^{-2(1/2)}\|\varphi\|.$
$\|u^{(2)}(\lambda)\|_{L^{2}(\Omega)}\leq\mathbf{c}|\Re\lambda|^{-1/2-2(1/2)}\|\varphi\|.$
$\|u^{(2)}(\lambda)\|_{L^{p^{\prime}}(\Omega)}\leq\mathbf{c}|\Re\lambda|^{-2(1/2)}\|\varphi\|.$
Next, by induction in $j$, $j\geq 1$, since by (2.15)
$(-\Delta+V-\lambda)u^{(j)}(\lambda)=ju^{(j-1)}(\lambda),\quad\gamma_{0}u^{(j)}=0,$
we argue as above to obtain
$\|u^{(j)}(\lambda)\|_{W^{2,q}(\Omega)}\leq\mathbf{c}_{j}|\Re\lambda|^{-j(1/2)}\|\varphi\|,$
$\|u^{(j)}(\lambda)\|_{L^{2}(\Omega)}\leq\mathbf{c}_{j}|\Re\lambda|^{-1/2-j(1/2)}\|\varphi\|,$
$\|u^{(j)}(\lambda)\|_{L^{p^{\prime}}(\Omega)}\leq\mathbf{c}_{j}|\Re\lambda|^{-j(1/2)}\|\varphi\|.$
These inequalities complete the proof. ∎
Next, we establish the following result.
Lemma 2.7.
Let $j\geq 0$ be an integer $(V_{1},V_{2})\in\mathscr{W}$, $\lambda\in\Pi_{2}$ and $\varphi\in\mathcal{B}$. Then we have
(2.48) $\|u_{V_{1}}^{(j)}(\lambda)(\varphi)-u_{V_{2}}^{(j)}(\lambda)(\varphi)\|_{W^{2,q}(\Omega)}\leq\mathbf{c}_{j}|\Re\lambda|^{-j/2}\|\varphi\|.$
(2.49) $\|u_{V_{1}}^{(j)}(\lambda)(\varphi)-u_{V_{2}}^{(j)}(\lambda)(\varphi)\|_{L^{2}(\Omega)}\leq\mathbf{c}_{j}|\Re\lambda|^{-(j+1)/2}\|\varphi\|.$
Proof.
Clearly, $u(\lambda):=u_{V_{1}}(\lambda)(\varphi)-u_{V_{2}}(\lambda)(\varphi)\in H_{0}^{1}(\Omega)$ satisfies
(2.50) $(-\Delta+V_{1}-\lambda)u(\lambda)=(V_{2}-V_{1})u_{V_{2}}(\lambda)(\varphi).$
That is $u(\lambda)=R_{V_{1}}(\lambda)((V_{2}-V_{1})u_{V_{2}}(\lambda)(\varphi))$. In light of (2.34), we obtain
$\|u(\lambda)\|_{W^{2,q}(\Omega)}$ $\leq\mathbf{c}\|(V_{2}-V_{1})u_{V_{2}}(\lambda)(\varphi)\|_{L^{2}(\Omega)}$
$\leq\mathbf{c}\|V_{2}-V_{1}\|_{L^{n}(\Omega)}\|u_{V_{2}}(\lambda)(\varphi)\|_{L^{p^{\prime}}(\Omega)}$
$\leq\mathbf{c}\|u_{V_{2}}(\lambda)(\varphi)\|_{L^{p^{\prime}}(\Omega)}.$
This and (2.40) yield
$\|u(\lambda)\|_{W^{2,q}(\Omega)}\leq\mathbf{c}\|\varphi\|.$
Similarly, we have
$\|u(\lambda)\|_{L^{2}(\Omega)}$ $\leq\mathbf{c}|\Re\lambda|^{-1/2}\|(V_{2}-V_{1})u_{V_{2}}(\lambda)(\varphi)\|_{L^{2}(\Omega)}$
$\leq\mathbf{c}|\Re\lambda|^{-1/2}\|u_{V_{2}}(\lambda)(\varphi)\|_{L^{p^{\prime}}(\Omega)}$
and then, again by (2.40),
$\|u(\lambda)\|_{L^{2}(\Omega)}\leq\mathbf{c}|\Re\lambda|^{-1/2}\|\varphi\|.$
By taking the derivative with respect to $\lambda$ of each side (2.50), we find
$(-\Delta+V_{1}-\lambda)u^{(1)}(\lambda)=u(\lambda)+(V_{2}-V_{1})u_{V_{2}}^{(1)}(\lambda)(\varphi).$
By proceeding as above, we obtain
$\|u^{(1)}(\lambda)\|_{W^{2,q}(\Omega)}\leq\mathbf{c}|\Re\lambda|^{-1/2}\|\varphi\|$
and
$\|u^{(1)}(\lambda)\|_{L^{2}(\Omega)}\leq\mathbf{c}|\Re\lambda|^{-1/2-1/2}\|\varphi\|.$
By using
$(-\Delta+V_{1}-\lambda)u^{(j)}(\lambda)=ju^{(j-1)}(\lambda)+(V_{2}-V_{1})u_{V_{2}}^{(j)}(\lambda)(\varphi),$
and an induction in $j$, $j\geq 1$, we get similarly as above
$\|u^{(j)}(\lambda)\|_{W^{2,q}(\Omega)}\leq\mathbf{c}_{j}|\Re\lambda|^{-j(1/2)}\|\varphi\|$
and
$\|u^{(j)}(\lambda)\|_{L^{2}(\Omega)}\leq\mathbf{c}_{j}|\Re\lambda|^{-(j+1)(1/2)}\|\varphi\|.$
The proof is then complete. ∎
We have from [13, Theorem 1.4.3.3] the following interpolation inequality
(2.51) $\|w\|_{W^{t,q}(\Omega)}\leq c\|w\|_{W^{2,q}(\Omega)}^{t/2}\|w\|_{L^{q}(\Omega)}^{1-t/2},$
where $c=c(n,\Omega,t)>0$ is a constant.
Recall that $\sigma=12/(2-t)$ and define
(2.52) $\Pi=\left\{\lambda=-\tau^{\sigma}+2i\tau\in\Pi_{2}\right\}.$
By combining (2.48), (2.49) and (2.51), and using that the trace map $w\in W^{t,q}(\Omega)\mapsto\gamma_{1}w\in L^{q}$ is bounded, we obtain
Corollary 2.2.
Let $j\geq 0$ be an integer, $(V_{1},V_{2})\in\mathscr{W}$, $\lambda=-\tau^{\sigma}+2i\tau\in\Pi$ and $\varphi\in\mathcal{B}$. Then we have
(2.53) $\|\Lambda_{V_{1}}^{(j)}(\lambda)(\varphi)-\Lambda_{V_{2}}^{(j)}(\lambda)(\varphi)\|_{L^{q}(\Gamma)}\leq\mathbf{\tilde{c}}\tau^{-3-j\sigma/2}\|\varphi\|,$
where $\mathbf{\tilde{c}}=\mathbf{\tilde{c}}(n,\Omega,V_{0},W_{0},t,j)$ is a constant.
3. Proof of Theorem 1.2
We recall for convenience that we have for all $(V_{1},V_{2})\in\mathscr{W}_{+}$
$\mathscr{D}_{+}(V_{1},V_{2}):=\sum_{k\geq 1}k^{-2/n}\left[|\lambda_{V_{1}}^{k}-\lambda_{V_{2}}^{k}|+\|\phi_{V_{2}}^{k}-\phi_{V_{1}}^{k}\|_{W^{2,q}(\Omega)}\right]<\infty,$
implying that
$\mathscr{D}(V_{1},V_{2}):=\sum_{k\geq 1}k^{-2/n}\left[|\lambda_{V_{1}}^{k}-\lambda_{V_{2}}^{k}|+\|\psi_{V_{2}}^{k}-\psi_{V_{1}}^{k}\|_{L^{q}(\Gamma)}\right]<\infty,$
Recall also that $\mathcal{B}:=W^{2-1/q,q}(\Gamma)\cap L^{q^{\prime}}(\Gamma)$ is equipped with the norm
$\|\varphi\|=\|\varphi\|_{W^{2-1/q,q}(\Gamma)}+\|\varphi\|_{L^{q^{\prime}}(\Gamma)},\quad\varphi\in\mathcal{B}.$
In the remaining part of this section, $\mathbf{c}=\mathbf{c}(n,\Omega,V_{0},W_{0},t)>0$ will denote a generic constant.
Before proceeding to the proof of Theorem 1.2, we will first prove the following intermediate result. Define
$\Sigma_{\tau^{\ast}}=\{\lambda_{\tau}=(\tau+i)^{2}\in\Sigma;\;\tau\geq\tau^{\ast}\},\quad\tau^{\ast}\geq 2.$
Proposition 3.1.
Let $(V_{1},V_{2})\in\mathscr{W}_{+}$, $\varphi\in\mathcal{B}$. There exists $\tau^{\ast}=\tau^{\ast}(n,\Omega,V_{0})$ such that for all $\lambda_{\tau}=(\tau+i)^{2}\in\Sigma_{\tau^{\ast}}$ we have
(3.1) $\|\Lambda_{V_{1}}(\lambda_{\tau})(\varphi)-\Lambda_{V_{1}}(\lambda_{\tau})(\varphi)\|\leq\mathbf{c}(\tau^{-3}+\tau^{\sigma+n+2}\mathscr{D}(V_{1},V_{2}))\|\varphi\|.$
Proof.
Define for $s\in[0,1]$ and $\tau\geq 2$
$\mu_{\tau}^{s}=-\tau^{\sigma}+s[\tau^{\sigma}+\tau^{2}-1]+2i\tau$
in such a way that $\mu_{\tau}^{0}=-\tau^{\sigma}+2i\tau\in\Pi$, where $\Pi$ is given by (2.52), and $\mu_{\tau}^{1}=\lambda_{\tau}=(\tau+i)^{2}\in\Sigma$.
Let $V_{1},V_{2}\in\mathscr{V}$. By applying Taylor’s formula, we get
(3.2) $u_{V_{j}}(\lambda_{\tau})=u_{V_{j}}(\mu_{\tau}^{0})(\varphi)+(\tau^{\sigma}+\tau^{2}-1)\int_{0}^{1}u_{V_{j}}^{(1)}(\mu_{\tau}^{s})(\varphi)ds,\quad j=1,2.$
Hence (2.53) implies
(3.3) $\mathbf{c}\|\Lambda_{V_{1}}(\lambda_{\tau})-\Lambda_{V_{2}}(\lambda_{\tau})\|_{L^{q}(\Gamma)}\leq\tau^{-3}\|\varphi\|$
$\hskip 85.35826pt+\tau^{\sigma}\int_{0}^{1}\|\Lambda_{V_{1}}(\mu_{\tau}^{s}(\varphi))-\Lambda_{V_{2}}(\mu_{\tau}^{s}(\varphi))\|_{L^{q}(\Gamma)}ds.$
In light of (2.5), we find an integer $\ell_{\ast}=\ell_{\ast}(n,\Omega,V_{0})>0$ and a constant $\mathbf{c}_{0}=(n,\Omega,V_{0})>0$ such that
(3.4) $\lambda_{V}^{k}\geq\mathbf{c}_{0}k^{2/n},\quad V\in\mathscr{V},\;k\geq\ell_{\ast}.$
On the other hand,
(3.5) $\Re\mu_{\tau}^{s}=\tau^{\sigma}-s[\tau^{\sigma}+\tau^{2}-1]\geq-\tau^{2},\quad s\in[0,1].$
Let $k_{\tau}$ be the smallest integer so that $k_{\tau}^{2/n}\geq 2\mathbf{c}_{0}^{-1}\tau^{2}$. We then choose $\tau^{\ast}=\tau^{\ast}(n,\Omega,V_{0})\geq 2$ so that $k_{\tau}\geq\ell_{\ast}$ for all $\tau\geq\tau^{\ast}$. Then, combining (3.4) and (3.5), we obtain
(3.6) $\lambda_{V}^{k}/2-\Re\mu_{\tau}^{s}\geq 0,\quad s\in[0,1],\;V\in\mathscr{V},\;k\geq k_{\tau}.$
For $V=V_{1}$ or $V=V_{2}$ and $\lambda\in\mathbb{C}\setminus\mathbb{R}$, we use the following decomposition $u_{V}(\lambda)(\varphi)=v_{V}(\lambda)(\varphi)+w_{V}(\lambda)(\varphi)$, where
$w_{V}(\lambda)(\varphi)=-\sum_{k<k_{\tau}}\frac{\langle\varphi,\psi_{V}^{k}\rangle}{\lambda_{V}^{k}-\lambda}\phi_{k},\quad v_{V}(\lambda)(g)=-\sum_{k\geq k_{\tau}}\frac{\langle\varphi,\psi_{V}^{k}\rangle}{\lambda_{V}^{k}-\lambda}\phi_{k}.$
We decompose
$v_{V_{1}}^{(1)}(\mu_{\tau}^{s})(\varphi)-v_{V_{2}}^{(1)}(\mu_{\tau}^{s})(\varphi)=$
$\hskip 71.13188pt\sum_{k\geq k_{\tau}}\frac{\langle\varphi,\psi_{V_{2}}^{k}\rangle}{(\lambda_{V_{2}}^{k}-\mu_{\tau}^{s})^{2}}\phi_{V_{2}}^{k}-\sum_{k\geq k_{\tau}}\frac{\langle\varphi,\phi_{V_{1}}^{k}\rangle}{(\lambda_{V_{1}}^{k}-\mu_{\tau}^{s})^{2}}\phi_{V_{1}}^{k}.$
in the form
(3.7) $v_{V_{1}}^{(1)}(\mu_{\tau}^{s})(\varphi)-v_{V_{2}}^{(1)}(\mu_{\tau}^{s})(\varphi)=\sum_{j=1}^{3}I_{j},$
where
$I_{1}=\sum_{k\geq k_{\tau}}\left(\frac{1}{(\lambda_{V_{2}}^{k}-\mu_{\tau}^{s})^{2}}-\frac{1}{(\lambda_{V_{1}}^{k}-\mu_{\tau}^{s})^{2}}\right)\langle\varphi,\psi_{V_{2}}^{k}\rangle\phi_{V_{2}}^{k},$
$I_{2}=\sum_{k\geq k_{\tau}}\frac{\langle\varphi,\psi_{V_{1}}^{k}-\psi_{V_{2}}^{k}\rangle}{(\lambda_{V_{1}}^{k}-\mu_{\tau}^{s})^{2}}\phi_{V_{2}}^{k},$
$I_{3}=\sum_{k\geq k_{\tau}}\frac{\langle\varphi,\psi_{V_{1}}^{k}\rangle}{(\lambda_{V_{1}}^{k}-\mu_{\tau}^{s})^{2}}(\phi_{V_{2}}^{k}-\phi_{V_{1}}^{k}).$
By using the formula
$\frac{1}{(\lambda_{V_{2}}^{k}-\mu_{\tau}^{s})^{2}}-\frac{1}{(\lambda_{V_{1}}^{k}-\mu_{\tau}^{s})^{2}}$
$\hskip 28.45274pt=\frac{(\lambda_{V_{1}}^{k}-\lambda_{V_{2}}^{k})}{(\lambda_{V_{1}}^{k}-\mu_{\tau}^{s})(\lambda_{V_{2}}^{k}-\mu_{\tau}^{s})}\left(\frac{1}{\lambda_{V_{2}}^{k}-\mu_{\tau}^{s}}+\frac{1}{\lambda_{V_{1}}^{k}-\mu_{\tau}^{s}}\right),$
we obtain
$\left|\frac{1}{(\lambda_{V_{2}}^{k}-\mu_{\tau}^{s})^{2}}-\frac{1}{(\lambda_{V_{1}}^{k}-\mu_{\tau}^{s})^{2}}\right|$ $\leq\frac{8|\lambda_{V_{1}}^{k}-\lambda_{V_{2}}^{k}|}{(\lambda_{V_{1}}^{k}\lambda_{V_{2}}^{k})^{2}}\left(\frac{1}{\lambda_{V_{2}}^{k}}+\frac{1}{\lambda_{V_{1}}^{k}}\right)$
$\leq\mathbf{c}k^{-6/n}|\lambda_{V_{1}}^{k}-\lambda_{V_{2}}^{k}|.$
By (2.11) and the continuity of the trace map $w\in W^{2,q}(\Omega)\mapsto\gamma_{1}w\in L^{q}(\Gamma)$, we have
(3.8) $\|\phi_{V_{2}}^{k}\|_{W^{2,q}(\Omega)}\leq\mathbf{c}(1+|\lambda_{V_{2}}^{k}|),\quad\|\psi_{V_{2}}^{k}\|_{L^{q}(\Gamma)}\leq\mathbf{c}(1+|\lambda_{V_{2}}^{k}|),\quad k\geq 1.$
Hence, the series
$\sum_{k\geq k_{\tau}}\left(\frac{1}{(\lambda_{V_{2}}^{k}-\mu_{\tau}^{s})^{2}}-\frac{1}{(\lambda_{V_{1}}^{k}-\mu_{\tau}^{s})^{2}}\right)\langle\varphi,\psi_{V_{2}}^{k}\rangle\phi_{V_{2}}^{k}$
converges in $W^{2,q}(\Omega)$ and
(3.9) $\|I_{1}\|_{W^{2,q}(\Omega)}\leq\mathbf{c}\sum_{k\geq k_{\tau}}k^{-2/n}|\lambda_{V_{1}}^{k}-\lambda_{V_{2}}^{k}|\|\varphi\|.$
We proceed similarly as above to obtain
(3.10) $\|I_{2}\|_{W^{2,q}(\Omega)}\leq\mathbf{c}\sum_{k\geq k_{\tau}}k^{-2/n}\|\psi_{V_{2}}^{k}-\psi_{V_{1}}^{k}\|_{L^{p}(\Gamma)}\|\varphi\|.$
Then (3.9) and (3.10) imply
(3.11) $\|\gamma_{1}I_{1}\|_{L^{q}(\Gamma)}+\|\gamma_{1}I_{2}\|_{L^{q}(\Gamma)}\leq\mathbf{c}\mathscr{D}(V_{1},V_{2})\|\varphi\|.$
On the other hand, since $\mathscr{D}_{+}(V_{1},V_{2})<\infty$, we verify that the series
$\sum_{k\geq k_{\tau}}\frac{\langle\varphi,\psi_{V_{1}}^{k}\rangle}{(\lambda_{V_{1}}^{k}-\mu_{\tau}^{s})^{2}}(\phi_{V_{2}}^{k}-\phi_{V_{1}}^{k})$
converges in $W^{2,q}(\Omega)$. Whence
$\gamma_{1}I_{3}=\sum_{k\geq k_{\tau}}\frac{\langle\varphi,\psi_{V_{1}}^{k}\rangle}{(\lambda_{V_{1}}^{k}-\mu_{\tau}^{s})^{2}}(\psi_{V_{2}}^{k}-\psi_{V_{1}}^{k}).$
As above, we derive from this identity
(3.12) $\|\gamma_{1}I_{3}\|_{L^{q}(\Gamma)}\leq\mathbf{c}\mathscr{D}(V_{1},V_{2})\|\varphi\|.$
In light of (3.7), a combination of (3.11) and (3.12) gives
(3.13) $\|\gamma_{1}v_{V_{1}}^{(1)}(\mu_{\tau}^{s})(\varphi)-\gamma_{1}v_{V_{2}}^{(1)}(\mu_{\tau}^{s})(\varphi)\|_{L^{q}(\Gamma)}\leq\mathbf{c}\mathscr{D}(V_{1},V_{2})\|\varphi\|,$
We proceed similarly to prove
$\mathbf{c}\|\gamma_{1}w_{V_{1}}^{(1)}(\mu_{\tau}^{s})(\varphi)-\gamma_{1}w_{V_{2}}^{(1)}(\mu_{\tau}^{s})(\varphi)\|_{L^{q}(\Gamma)}\leq\frac{1}{\tau^{3}}\sum_{k<k_{\tau}}(1+|\lambda_{V_{2}}^{k}|)^{2}|\lambda_{V_{1}}^{k}-\lambda_{V_{2}}^{k}|\|\varphi\|$
$\hskip 113.81102pt+\frac{1}{\tau^{2}}\sum_{k<k_{\tau}}(1+|\lambda_{V_{1}}^{k}|+|\lambda_{V_{2}}^{k}|)\|\psi_{V_{1}}^{k}-\psi_{V_{2}}^{k}\|_{L^{q}(\Gamma)}\|\varphi\|.$
In light of the definition of $k_{\tau}$, we verify that
(3.14) $\|\gamma_{1}w_{V_{1}}^{(1)}(\mu_{\tau}^{s})(\varphi)-\gamma_{1}w_{V_{2}}^{(1)}(\mu_{\tau}^{s})(\varphi)\|_{L^{q}(\Gamma)}\leq\mathbf{c}\tau^{n+2}\mathscr{D}(V_{1},V_{2})\|\varphi\|$
Using (3.13) and (3.14) in (3.3), we obtain
$\|\Lambda_{V_{1}}(\lambda_{\tau})-\Lambda_{V_{2}}(\lambda_{\tau})\|_{L^{q}(\Gamma)}\leq\mathbf{c}(\tau^{-3}+\tau^{\sigma+n+2}\mathscr{D}(V_{1},V_{2}))\|\varphi\|.$
This is the expected inequality. ∎
Remark 3.1.
In the previous proof we only use Corollary 2.2 with $j=0$. Corollary 2.2 also serves to explain why it is not possible, contrary to what we did in [11], to use the higher-order Taylor formula to remove the condition $\mathscr{D}_{+}(V_{1},V_{2})<\infty$ in Theorem 1.2. For example, if we use the Taylor formula of order $1$, then in light of (2.53), we obtain $(\tau^{-3}+\tau^{-3+\sigma/2})\|\varphi\|$ instead of $\tau^{-3}\|\varphi\|$. We can then see in the remaining part of the proof of Theorem 1.2 below that this is not sufficient to control this term by a negative power of $\tau$ because we use functions $\varphi$ satisfying $\|\varphi\|=O(\tau^{2})$.
We are now ready to complete the proof of Theorem 1.2. To this end, let $\varphi_{\lambda,\omega}(x)=e^{i\sqrt{\lambda}\omega\cdot x}$, $\lambda\in\mathbb{C}\setminus(-\infty,0]$, with the standard choice of the branch of the square root, and $\omega\in\mathbb{S}^{n-1}$. Define
$S_{V}(\lambda,\omega,\theta)=\int_{\Gamma}\Lambda_{V}(\lambda)(\varphi_{\lambda,\omega})\varphi_{\lambda,-\theta}d\sigma,\quad\lambda\in\rho(A_{V})\setminus(-\infty,0],\;\omega,\;\theta\in\mathbb{S}^{n-1}.$
Following [14, Lemma 2.2] we have for all $\lambda\in\rho(A_{V})\setminus(-\infty,0]$ and $\omega,\;\theta\in\mathbb{S}^{n-1}$
(3.15) $S_{V}(\lambda,\omega,\theta)=-\frac{\lambda}{2}|\theta-\omega|^{2}\int_{\Omega}e^{-i\sqrt{\lambda}(\theta-\omega)\cdot x}dx$
$\hskip 85.35826pt+\int_{\Omega}e^{-i\sqrt{\lambda}(\theta-\omega)\cdot x}Vdx-\int_{\Omega}R_{V}(\lambda)(V\varphi_{\lambda,\omega})V\varphi_{\lambda,-\theta}dx.$
We fix $\xi\in\mathbb{R}^{n}$ and $\eta\in\mathbb{S}^{n-1}$ such that $\eta\bot\xi$. For $\tau>1$, let
$\left\{\begin{array}[]{ll}&\theta_{\tau}=c_{\tau}\eta+(2\tau)^{-1}\xi,\\ &\omega_{\tau}=c_{\tau}\eta-(2\tau)^{-1}\xi,\\ &\sqrt{\lambda_{\tau}}=\tau+i.\end{array}\right$.
Let $(V_{1},V_{2})\in\mathscr{W}_{+}$ and set $V=(V_{1}-V_{2})\chi_{\Omega}$ and $\lambda_{\tau}=(\tau+i)^{2}\in\Sigma_{\tau^{\ast}}$, where $\Sigma_{\tau^{\ast}}$ is as in Proposition 3.1. By using the formula (3.15), we get
(3.16) $|\widehat{V}((1+i/\tau)\xi)|\leq|S_{V_{1}}(\lambda_{\tau},\omega_{\tau},\theta_{\tau})-S_{V_{2}}(\lambda_{\tau},\omega_{\tau},\theta_{\tau})|$
$\quad+\left|\int_{\Omega}R_{V_{1}}(\lambda)(V_{1}\varphi_{\lambda,\omega})V_{1}\varphi_{\lambda_{\tau},-\theta_{\tau}}dx\right|+\left|\int_{\Omega}R_{V_{2}}(\lambda)(V_{2}\varphi_{\lambda_{\tau},\omega_{\tau}})V_{2}\varphi_{\lambda,-\theta}dx\right|.$
Here $\widehat{V}$ denotes the Laplace-Fourier transform of $V$.
Let $r=p_{1/4}=4n/(2n+1)$ and $r^{\prime}=p^{\prime}_{1/4}=4n/(2n-1)$ ($p_{\theta}$ and $p^{\prime}_{\theta}$ are defined in (2.25)). For $W=V_{1},V_{2}$ and $\zeta=\omega_{\tau},-\theta_{\tau}$, Hölder’s inequality yields
$\|W\varphi_{\lambda_{\tau},\zeta}\|_{L^{r}(\Omega)}\leq\|W\|_{L^{n/2}(\Omega)}\|\varphi_{\lambda_{\tau},\zeta}\|_{L^{\tilde{q}}(\Omega)},$
where $\tilde{q}=4n/(2n-7)$. Thus, we have
$\|W\varphi_{\lambda_{\tau},\zeta}\|_{L^{r}(\Omega)}\leq\mathbf{c}.$
which, in combination with (2.26), gives for $j=1,2$
(3.17) $\left|\int_{\Omega}R_{V_{j}}(\lambda)(V_{j}\varphi_{\lambda_{\tau},\omega_{\tau}})V_{j}\varphi_{\lambda,-\theta}dx\right|\leq\mathbf{c}\tau^{-1/2}.$
Using (3.17) in (3.16), we obtain
(3.18) $|\widehat{V}((1+i/\tau)\xi)|\leq|S_{V_{1}}(\lambda_{\tau},\omega_{\tau},\theta_{\tau})-S_{V_{2}}(\lambda_{\tau},\omega_{\tau},\theta_{\tau})|+\mathbf{c}\tau^{-1/2}.$
According to the mean-value theorem, we have
$|\widehat{V}(\xi)|\leq|\widehat{V}((1+i/\tau)\xi)|+\tau^{-1}\sup_{0\leq s\leq 1}|\nabla\widehat{V}((1+is/\tau)\xi)\cdot\xi|.$
But we have for $1\leq k\leq n$
$|\partial_{k}\widehat{V}((1+i/\tau)\xi)((1+is/\tau)\xi)|=|\widehat{x_{k}V}((1+is/\tau)\xi)|\leq e^{\mathbf{c}\tau^{-1}|\xi|}.$
Whence
$|\widehat{V}(\xi)|\leq|\hat{V}((1+i/\tau)\xi)|+\mathbf{c}\tau^{-1}|\xi|e^{\mathbf{c}\tau^{-1}|\xi|}.$
This in (3.18) yields
(3.19) $|\widehat{V}(\xi)|\leq|S_{V_{1}}(\lambda_{\tau},\omega_{\tau},\theta_{\tau})-S_{V_{2}}(\lambda_{\tau},\omega_{\tau},\theta_{\tau})|+\mathbf{c}\tau^{-1/2}+e^{\mathbf{c}\tau^{-1}|\xi|}.$
On the other hand, we have
$|S_{V_{1}}(\lambda_{\tau},\omega_{\tau},\theta_{\tau})-S_{V_{2}}(\lambda_{\tau},\omega_{\tau},\theta_{\tau})|$
$\hskip 56.9055pt\leq\|\Lambda_{V_{1}}(\lambda_{\tau})(\varphi_{\lambda,\omega})-\Lambda_{V_{2}}(\lambda_{\tau})(\varphi_{\lambda,\omega})\|_{L^{q}(\Gamma)}\|\varphi_{\lambda,-\theta}\|_{L^{q^{\prime}}(\Gamma)}.$
$\hskip 56.9055pt\leq\mathbf{c}\|\Lambda_{V_{1}}(\lambda_{\tau})(\varphi_{\lambda,\omega})-\Lambda_{V_{2}}(\lambda_{\tau})(\varphi_{\lambda,\omega})\|_{L^{q}(\Gamma)},$
which, in light of (3.1), yields
$\mathbf{c}|S_{V_{1}}(\lambda_{\tau},\omega_{\tau},\theta_{\tau})-S_{V_{2}}(\lambda_{\tau},\omega_{\tau},\theta_{\tau})|\leq\tau^{-1/2}+\tau^{\sigma+n+4}\mathscr{D}(V_{1},V_{2}),$
where we used
$\|\varphi_{\lambda,\omega}\|\leq\mathbf{c}\tau^{2}.$
Therefore, we obtain from (3.19)
(3.20) $\int_{|\xi|\leq\tau^{\alpha}}(1+|\xi|^{2})^{-1}|\widehat{V}(\xi)|^{2}d\xi\leq\mathbf{c}\tau^{2(\sigma+n+4)+n\alpha}\mathscr{D}(V_{1},V_{2})^{2}$
$\hskip 156.49014pt+\mathbf{c}\tau^{-1+n\alpha}+\mathbf{c}\tau^{n\alpha}e^{\mathbf{c}\tau^{-1+\alpha}},$
where $\alpha>0$ will be chosen later.
We verify that
(3.21) $\int_{|\xi|>\tau^{\alpha}}(1+|\xi|^{2})^{-1}|\widehat{V}(\xi)|^{2}d\xi\leq\tau^{-2\alpha}\int_{\mathbb{R}^{n}}|\widehat{V}(\xi)|^{2}dx.$
According to Parseval’s identity, we have
$\int_{\mathbb{R}^{n}}|\widehat{V}(\xi)|^{2}dx=\int_{\mathbb{R}^{n}}|V(\xi)|^{2}dx=\|V_{1}-V_{2}\|_{L^{2}(\Omega)}\leq\mathbf{c}$
and then (3.21) implies
(3.22) $\int_{|\xi|>\tau^{\alpha}}(1+|\xi|^{2})^{-1}|\widehat{V}(\xi)|^{2}d\xi\leq\mathbf{c}\tau^{-2\alpha}.$
Putting together (3.20) and (3.22), we get
(3.23) $\|V\|_{H^{-1}(\mathbb{R}^{n})}^{2}\leq\mathbf{c}\tau^{2(\sigma+n+4)+n\alpha}\mathscr{D}(V_{1},V_{2})^{2}$
$\hskip 128.0374pt+\mathbf{c}\tau^{-1+n\alpha}+\mathbf{c}\tau^{n\alpha}e^{\mathbf{c}\tau^{-1+\alpha}}+\mathbf{c}\tau^{-2\alpha}.$
Assume first that $0<\alpha<1$. In this case since
$\mathbf{c}_{\alpha}:=\sup_{\tau\geq 1}\{\mathbf{c}\tau^{(n+2)\alpha}e^{\mathbf{c}\tau^{-1+\alpha}}\}<\infty,$
(3.23) implies
(3.24) $\|V\|_{H^{-1}(\mathbb{R}^{n})}^{2}\leq\mathbf{c}\tau^{2(\sigma+n+4)+n\alpha}\mathscr{D}(V_{1},V_{2})^{2}$
$\hskip 128.0374pt+\mathbf{c}\tau^{-1+n\alpha}+(\mathbf{c}+\mathbf{c}_{\alpha})\tau^{-2\alpha}.$
By taking $\alpha=1/(n+2)$ in (3.24), we obtain
$\|V\|_{H^{-1}(\mathbb{R}^{n})}\leq\mathbf{c}\tau^{\sigma+n+4+n/(2n+4)}\mathscr{D}(V_{1},V_{2})+\mathbf{c}\tau^{-1/(n+2)}.$
This and the fact that $\|V_{1}-V_{2}\|_{H^{-1}(\Omega)}\leq\|V\|_{H^{-1}(\mathbb{R}^{n})}$ yield
(3.25) $\|V_{1}-V_{2}\|_{H^{-1}(\Omega)}\leq\mathbf{c}\tau^{\sigma+n+4+n/(2n+4)}\mathscr{D}(V_{1},V_{2})+\mathbf{c}\tau^{-1/(n+2)}.$
By minimizing with respect to $\tau$, we obtain (1.1) from (3.25).
4. Extension to the anisotropic case
We briefly explain in this section how to modify the previous proof to obtain an extension of Theorem 1.2 to the anisotropic case. We reuse the definitions and notations from the previous sections.
Suppose that $\Omega$ is $C^{\infty}$ smooth. Let $\mathbf{g}=(g_{k\ell})$ be a Riemannian metric $C^{\infty}$ in $\overline{\Omega}$. We assume that $\mathbf{g}$ is simple in $\Omega$, which means that $\Gamma$ is strictly convex with respect to the metric $\mathbf{g}$ and for all $x\in\overline{\Omega}$ the exponential map $\exp_{x}:\exp_{x}^{-1}(\overline{\Omega})\rightarrow\overline{\Omega}$ is a diffeomorphism. We note that $\mathbf{g}$ can be extended to a simple metric in $\Omega_{1}\Supset\Omega$.
All the results of the previous sections up to Proposition 3.1 remain valid. We need to replace the functions $\varphi_{\lambda,\omega}$ by geometric optic solutions associated with the metric $\mathbf{g}$.
In the following, $\Delta$ represents the Laplace-Beltrami operator associated with $\mathbf{g}$. Since $\mathbf{g}$ is simple, there is $\phi\in C^{\infty}(\overline{\Omega})$ a solution to the eikonal equation
(4.1) $|\nabla\phi|^{2}:=\sum_{k,\ell=1}^{n}g^{k\ell}\frac{\partial\phi}{\partial x_{k}}\frac{\partial\phi}{\partial x_{\ell}}=1.$
Here $(g^{k\ell}):=\mathbf{g}^{-1}$ is the inverse of the metric $\mathbf{g}$. Moreover, the transport equation
(4.2) $\sum_{k,\ell=1}^{n}g^{k\ell}\frac{\partial\phi}{\partial x_{k}}\frac{\partial a}{\partial x_{\ell}}+\frac{1}{2}(\Delta\phi)a=0$
admits a solution $a\in H^{3}(\Omega)$. These statements are obtained by slightly modifying the proof in the second part of [5, Section 4]. The function $\phi$ is generally called a phase and the function $a$ an amplitude.
In the following, $\phi$ will denote a fixed solution of the eikonal equation (4.1) and $a,b\in H^{3}(M)$ will denote two solutions of the transport equation (4.2).
Let $A_{0}$ be the unbounded operator acting on $L^{2}(\Omega)$ as follows
$A_{0}u=-\Delta u,\quad u\in D(A_{0}):=H^{2}(\Omega)\cap H_{0}^{1}(\Omega).$
For $\zeta\in\{a,b\}$ and $\lambda_{\tau}=(\tau+i)^{2}\in\Sigma$, let
$\varphi^{\pm}_{\zeta,\tau}:=e^{\pm\sqrt{\lambda_{\tau}}\phi}\zeta-(A_{0}-\lambda_{\tau})^{-1}(e^{\pm\sqrt{\lambda_{\tau}}\phi}\Delta\zeta).$
We verify that $\varphi^{\pm}_{\zeta,\tau}\in H^{2}(\Omega)$ and $(\Delta+\lambda_{\tau})\varphi^{\pm}_{\zeta,\tau}=0$.
In the following, $\mathbf{c}_{0}=\mathbf{c}_{0}(n,\Omega)>0$ will denote a generic constant.
If $\lambda_{\tau}\in\Sigma\subset\Pi_{0}$, then we have from (2.17)
$\|(A_{0}-\lambda_{\tau})^{-1}(e^{\pm\sqrt{\lambda_{\tau}}\phi}\Delta\zeta)\|_{L^{p^{\prime}}(\Omega)}\leq\mathbf{c}_{0}\|\zeta\|_{H^{2}(\Omega)}.$
Whence
(4.3) $\|\varphi^{\pm}_{\zeta,\tau}\|_{L^{p^{\prime}}(\Omega)}\leq\mathbf{c}_{0}\|\zeta\|_{H^{2}(\Omega)}.$
Furthermore, as $\varphi^{\pm}_{\zeta,\tau}=e^{\pm\sqrt{\lambda_{\tau}}\phi}\zeta$ on $\Gamma$, we obtain
$\|\varphi^{\pm}_{\zeta,\tau}\|_{L^{q^{\prime}}(\Gamma)}\leq\mathbf{c}_{0}\|\zeta\|_{L^{q^{\prime}}(\Gamma)}.$
This inequality, combined with the fact that $H^{3}(\Omega)$ is continuously embedded in $W^{1,q^{\prime}}(\Omega)$ ([12, (7.30)] shows that $H^{2}(\Omega)$ in continuously embedded in $L^{q^{\prime}}(\Omega)$) and the continuity of the trace map $w\in W^{1,q^{\prime}}(\Omega)\mapsto w_{|\Gamma}\in L^{q^{\prime}}(\Gamma)$, yields
(4.4) $\|\varphi^{\pm}_{\zeta,\tau}\|_{L^{q^{\prime}}(\Gamma)}\leq\mathbf{c}_{0}\|\zeta\|_{H^{3}(\Omega)}.$
We use (4.4) to obtain
(4.5) $\|\varphi^{\pm}_{\zeta,\tau}\|\leq\mathbf{c}_{0}\tau^{2}\|\zeta\|_{H^{3}(\Omega)}.$
For $n\geq 7$, let $s=n/2$ and $\mathscr{V}_{+}=\mathscr{V}$ and, for $n=5,6$, let $s>4n/(n+2)$ and set
$\mathscr{V}_{+}:=\{V\in\mathscr{V}\cap L^{s}(\Omega);\;\|V\|_{L^{s}(\Omega)}\leq\varpi\}.$
We assume that $\varpi>0$ is chosen sufficiently large in such a way to guarantee that $\mathscr{V}_{+}\neq\emptyset$. Also, let
$\theta=\theta(n,s):=\frac{sn^{2}+2sn-4n^{2}}{s(2n+4)}\in(0,1/2).$
Let $V\in\mathscr{V}_{+}$, $r=2n/(n+2\theta)$. Applying Hölder’s inequality, we obtain from (4.3)
$\|V\varphi^{\pm}_{\zeta,\tau}\|_{L^{r}(\Omega)}\leq\|V\|_{L^{s}(\Omega)}\|\varphi^{\pm}_{\zeta,\tau}\|_{L^{p^{\prime}}(\Omega)}\leq\mathbf{c}\|\zeta\|_{H^{2}(\Omega)}.$
Here and in the remaining part of this section, $\mathbf{c}>0$ will denote a generic constant depending only on $n$, $\Omega$, $\mathbf{g}$, $V_{0}$, $s$ and $\varpi$.
Let $d\mu$ denotes the Riemannian measure on $\Omega$. From (2.26) and (4.3), we obtain
(4.6) $\left|\int_{\Omega}R_{V}(\lambda)(V\varphi^{\pm}_{a,\tau})V\varphi^{\pm}_{b,\tau}d\mu\right|\leq\mathbf{c}\tau^{-\vartheta}\|a\|_{H^{2}(\Omega)}\|b\|_{H^{2}(\Omega)},$
where $\vartheta=1-2\theta$.
Let $V_{1},V_{2}\in\mathscr{V}_{+}$ and set
$\varphi_{\tau}^{+}:=\varphi^{+}_{a,\tau},\quad\varphi_{\tau}^{-}:=\varphi^{-}_{b,\tau}.$
As in the isotropic case, we establish the following formula
$\int_{\Omega}(V_{1}-V_{2})\varphi_{\tau}^{+}\varphi_{\tau}^{-}d\sigma=\int_{\Gamma}(\Lambda_{V_{1}}(\lambda_{\tau})-\Lambda_{V_{2}}(\lambda_{\tau}))(\varphi_{\tau}^{+})\varphi_{\tau}^{-}d\sigma$
$\hskip 85.35826pt+\int_{\Omega}R_{V_{1}}(V_{1}\varphi_{\tau}^{+})V_{1}\varphi_{\tau}^{-}d\mu-\int_{\Omega}R_{V_{2}}(V_{2}\varphi_{\tau}^{+})V_{2}\varphi_{\tau}^{-}d\mu.$
Then (4.6) implies
$\left|\int_{\Omega}(V_{1}-V_{2})\varphi_{\tau}^{+}\varphi_{\tau}^{-}d\sigma\right|\leq\|(\Lambda_{V_{1}}(\lambda_{\tau})-\Lambda_{V_{2}}(\lambda_{\tau}))(\varphi_{\tau}^{+})\|_{L^{q}(\Gamma)}\|\varphi_{\tau}^{-}\|_{L^{q^{\prime}}(\Gamma)}$
$\hskip 213.39566pt+\mathbf{c}\tau^{-\vartheta}\|a\|_{H^{2}(\Omega)}\|b\|_{H^{2}(\Omega)},$
which, combined with (4.4), implies
(4.7) $\left|\int_{\Omega}(V_{1}-V_{2})\varphi_{\tau}^{+}\varphi_{\tau}^{-}d\sigma\right|\leq\mathbf{c}\|(\Lambda_{V_{1}}(\lambda_{\tau})-\Lambda_{V_{2}}(\lambda_{\tau}))(\varphi_{\tau}^{+})\|_{L^{q}(\Gamma)}\|b\|_{H^{3}(\Omega)}$
$\hskip 199.16928pt+\mathbf{c}\tau^{-\vartheta}\|a\|_{H^{2}(\Omega)}\|b\|_{H^{2}(\Omega)}.$
Now, the generic constant $\mathbf{c}$ can also depend on $W_{0}$.
Assume that $(V_{1},V_{2})\in(\mathscr{V}_{+}\times\mathscr{V}_{+})\cap\mathscr{W}_{+}$. Then we obtain from (3.1)
$\mathbf{c}\|\Lambda_{V_{1}}(\lambda_{\tau})(\varphi_{\tau}^{+})-\Lambda_{V_{1}}(\lambda_{\tau})(\varphi_{\tau}^{+})\|\leq(\tau^{-3}+\tau^{\sigma+n+2}\mathscr{D}(V_{1},V_{2}))\|\varphi_{\tau}^{+}\|.$
This and (4.5) give
(4.8) $\mathbf{c}\|\Lambda_{V_{1}}(\lambda_{\tau})(\varphi_{\tau}^{+})-\Lambda_{V_{1}}(\lambda_{\tau})(\varphi_{\tau}^{+})\|\leq(\tau^{-1}+\tau^{\sigma+n+4}\mathscr{D}(V_{1},V_{2}))\|a\|_{H^{3}(\Omega)}.$
Combining (4.7) and (4.8), we find
(4.9) $\mathbf{c}\left|\int_{\Omega}(V_{1}-V_{2})\varphi_{\tau}^{+}\varphi_{\tau}^{-}d\sigma\right|\leq(\tau^{-1}+\tau^{\sigma+n+4}\mathscr{D}(V_{1},V_{2}))\|b\|_{H^{3}(\Omega)}\|a\|_{H^{3}(\Omega)}$
$\hskip 199.16928pt+\tau^{-\vartheta}\|a\|_{H^{2}(M)}\|b\|_{H^{2}(\Omega)}.$
On the other hand, using $L^{2}$ resolvent estimate, we get
$\left|\int_{\Omega}(V_{1}-V_{2})abd\sigma\right|\leq\left|\int_{\Omega}(V_{1}-V_{2})\varphi_{\tau}^{+}\varphi_{\tau}^{-}d\sigma\right|+\mathbf{c}\tau^{-1}\|a\|_{H^{2}(\Omega)}\|b\|_{H^{2}(\Omega)}.$
This inequality in (4.9) gives
(4.10) $\mathbf{c}\left|\int_{\Omega}(V_{1}-V_{2})abd\sigma\right|\leq(\tau^{-1}+\tau^{\sigma+n+4}\mathscr{D}(V_{1},V_{2}))\|b\|_{H^{3}(\Omega)}\|a\|_{H^{3}(\Omega)}$
$\hskip 199.16928pt+\tau^{-\vartheta}\|a\|_{H^{2}(M)}\|b\|_{H^{2}(\Omega)}.$
Let $\mathbf{I}$ be the geodesic ray transform and $\mathbf{N}=\mathbf{I}^{\ast}\mathbf{I}:L^{2}(\Omega)\rightarrow H^{1}(\Omega_{1})$. The properties of the operators $\mathbf{I}$ and $\mathbf{N}$ that we will use below are borrowed from [5, Section 2].
Under the assumption that $V_{1}-V_{2}\in H^{2}(\Omega)$ and
$\|V_{1}-V_{2}\|_{H^{2}(\Omega)}\leq\varpi,$
proceeding as in [5, Section 5], we obtain from (4.10)
(4.11) $\mathbf{c}\|\mathbf{N}(V_{1}-V_{2})\|_{L^{2}(\Omega_{1})}^{2}\leq\tau^{\sigma+n+4}\mathscr{D}(V_{1},V_{2})+\tau^{-\vartheta}.$
where we used that
$\|\mathbf{N}(V_{1}-V_{2})\|_{H^{j+1}(\Omega_{1})}\leq\mathbf{c}\|V_{1}-V_{2}\|_{H^{j}(\Omega)},\quad j=0,1.$
On the other hand, we have
$\|V_{1}-V_{2}\|_{L^{2}(\Omega)}\leq$ $\mathbf{c}\|\mathbf{N}(V_{1}-V_{2})\|_{H^{1}(\Omega_{1})}$
$\leq\|\mathbf{N}(V_{1}-V_{2})\|_{L^{2}(\Omega_{1})}^{1/2}\|\mathbf{N}(V_{1}-V_{2})\|_{H^{2}(\Omega_{1})}^{1/2}$
$\leq\mathbf{c}\|\mathbf{N}(V_{1}-V_{2})\|_{L^{2}(\Omega_{1})}^{1/2}.$
This and (4.11) imply
(4.12) $\mathbf{c}\|V_{1}-V_{2}\|_{L^{2}(\Omega_{1})}^{4}\leq\tau^{\sigma+n+4}\mathscr{D}(V_{1},V_{2})+\tau^{-\vartheta}.$
Define
$\tilde{\mathscr{W}}_{+}=\{(V_{1},V_{2})\in(\mathscr{V}_{+}\times\mathscr{V}_{+})\cap\mathscr{W}_{+};\;V_{1}-V_{2}\in H^{2}(\Omega),\,\|V_{1}-V_{2}\|_{H^{2}(\Omega)}\leq\varpi\}.$
Minimizing the right hand side of (4.12) with respect to $\tau$, we obtain the following result, where
$\tilde{\beta}:=\frac{\vartheta}{4\sigma+4n+16+4\vartheta}.$
Theorem 4.1.
For all $(V_{1},V_{2})\in\tilde{\mathscr{W}}_{+}$ we have
$\|V_{1}-V_{2}\|_{L^{2}(\Omega)}\leq\mathbf{c}\mathscr{D}(V_{1},V_{2})^{\tilde{\beta}}$.
Appendix A Coercivity of the sesquilinear form $\mathfrak{a}_{V}$
Before proving the coercivity of $\mathfrak{a}_{V}$, we establish the following preliminary lemma. We recall for convenience that
$\mathbf{e}=\sup\left\{\|w\|_{L^{p^{\prime}}(\Omega)};\;u\in H^{1}(\Omega),\;\|u\|_{H^{1}(\Omega)}=1\right\}.$
$\mathbf{p}=\sup\left\{\|u\|_{H^{1}(\Omega)};\;u\in H_{0}^{1}(\Omega),\;\|\nabla u\|_{L^{2}(\Omega)}=1\right\}.$
Lemma A.1.
Let $V\in L^{n/2}(\Omega)$ and denote by $\mathfrak{b}_{V}$ the sesquilinear form defined by
$\mathfrak{b}_{V}(u,v)=\int_{\Omega}Vu\overline{v}dx,\quad u,v\in H^{1}(\Omega).$
Then $\mathfrak{b}$ is bounded with
(A.1) $|\mathfrak{b}_{V}(u,v)|\leq\mathbf{e}^{2}\|V\|_{L^{n/2}(\Omega)}\|u\|_{H^{1}(\Omega)}\|v\|_{H^{1}(\Omega)},\quad u,v\in H^{1}(\Omega).$
Moreover, for all $u\in H^{1}(\Omega)$, $\ell_{V}(u):v\in H_{0}^{1}(\Omega)\mapsto\mathfrak{b}_{V}(u,v)$ belongs to $H^{-1}(\Omega)$ with
(A.2) $\|\ell_{V}(u)\|_{H^{-1}(\Omega)}\leq\mathbf{e}^{2}\|V\|_{L^{n/2}(\Omega)}\|u\|_{H^{1}(\Omega)}.$
Proof.
Let $u,v\in H^{1}(\Omega)$. By applying Cauchy-Schwarz’s inequality, we get
(A.3) $|\mathfrak{b}_{V}(u,v)|\leq\|Vu^{2}\|_{L^{1}(\Omega)}^{1/2}\|Vv^{2}\|_{L^{1}(\Omega)}^{1/2}.$
It follows from Hölder’s inequality
(A.4) $\|Vw^{2}\|_{L^{1}(\Omega)}^{1/2}\leq\|V\|_{L^{n/2}(\Omega)}^{1/2}\|w\|_{L^{p^{\prime}}(\Omega)}\leq\mathbf{e}\|V\|_{L^{n/2}(\Omega)}^{1/2}\|w\|_{H^{1}(\Omega)},\quad w\in\{u,v\}.$
Then (A.4) in (A.3) yields (A.1).
Let $u\in H^{1}(\Omega)$. From (A.1), we obtain
$|\ell_{V}(u)(v)|\leq\mathbf{e}^{2}\|V\|_{L^{n/2}(\Omega)}\|u\|_{H^{1}(\Omega)}\|v\|_{H_{0}^{1}(\Omega)},\quad v\in H_{0}^{1}(\Omega).$
Therefore, $\ell_{V}(u)\in H^{-1}(\Omega)$ and (A.2) holds. ∎
Let $V\in L^{n/2}(\Omega,\mathbb{R})$ and recall that $\mathfrak{a}_{V}$ is given by
$\mathfrak{a}_{V}(u,v)$ $=\int_{\Omega}\left(\nabla u\cdot\nabla\overline{v}+Vu\overline{v}\right)dx$
$=\int_{\Omega}\nabla u\cdot\nabla\overline{v}dx+\mathfrak{b}_{V}(u,v),\quad u,v\in H_{0}^{1}(\Omega).$
In light of Lemma A.1, $\mathfrak{a}_{V}$ is bounded with
$|\mathfrak{a}_{V}(u,v)|\leq\left(1+\|V\|_{L^{n/2}(\Omega)}\right)\|u\|_{H^{1}(\Omega)}\|v\|_{H^{1}(\Omega)}.$
As $C_{0}^{\infty}(\Omega)$ is dense in $L^{n/2}(\Omega)$, we find $\tilde{V}\in C_{0}^{\infty}(\Omega)$ so that
$\|V-\tilde{V}\|_{L^{n/2}(\Omega)}\leq\frac{1}{2\mathbf{p}^{2}\mathbf{e}^{2}}.$
Therefore, we have for $u\in H_{0}^{1}(\Omega)$
$\mathfrak{a}_{V}(u,u)$ $\geq\mathbf{p}^{-2}\|u\|_{H^{1}(\Omega)}^{2}-\|\tilde{V}\|_{L^{\infty}(\Omega)}\|u\|_{L^{2}(\Omega)}^{2}$
$\hskip 85.35826pt-\|V-\tilde{V}\|_{L^{n/2}(\Omega)}\|u\|_{L^{p^{\prime}}(\Omega)}^{2}$
$\geq\frac{\mathbf{p}^{-2}}{2}\|u\|_{H^{1}(\Omega)}^{2}-\|\tilde{V}\|_{L^{\infty}(\Omega)}\|u\|_{L^{2}(\Omega)}^{2}.$
This shows that $\mathfrak{a}_{V}$ is coercive.
Appendix B BVPs with unbounded lower order term
We consider the BVP
(B.1) $\left\{\begin{array}[]{ll}(-\Delta+V-\lambda)u=f\quad&\mbox{in}\;\Omega,\\ u=\varphi&\mbox{on}\;\Gamma.\end{array}\right$.
Theorem B.1.
Let $V\in L^{n/2}(\Omega,\mathbb{R})$ and $\lambda\in\rho(A_{V})$. For all $f\in L^{q}(\Omega)$ and $\varphi\in W^{2-1/q,q}(\Gamma)$ the BVP (B.1) has a unique solution $u=u_{V}(\lambda)(\varphi,f)\in W^{2,q}(\Omega)$. In addition, there exists a constant $\varkappa=\varkappa(n,\Omega,V,\lambda)>0$ such that
(B.2) $\|u_{V}(\lambda)(\varphi,f)\|_{W^{2,q}(\Omega)}\leq\varkappa\left(\|f\|_{L^{q}(\Omega)}+\|\varphi\|_{W^{2-1/q,q}(\Gamma)}\right).$
Proof.
In this proof, $\mathbf{c}_{0}=\mathbf{c}_{0}(n,\Omega)>0$ and $\mathbf{c}=\mathbf{c}(n,\Omega,V_{0})>0$ are generic constants.
Let $F\in W^{2,q}(\Omega)$ so that $\gamma_{0}F=\varphi$ and
$\|F\|_{W^{2,q}(\Omega)}\leq\mathbf{c}_{0}\|\varphi\|_{W^{2-1/q,q}(\Gamma)}.$
If $G=-(-\Delta+V-\lambda)F+f$ then $G\in L^{q}(\Omega)$ and
(B.3) $\|G\|_{L^{q}(\Omega)}\leq\mathbf{c}\left(\|f\|_{L^{q}(\Omega)}+\|\varphi\|_{W^{2-1/q,q}(\Gamma)}\right).$
By density, we can find a sequence $(G_{k})$ in $L^{2}(\Omega)$ converging in $L^{q}(\Omega)$ to $G$. Set for any $k$
(B.4) $v_{k}=R_{V}(\lambda)G_{k}.$
Then $v_{k}\in H_{0}^{1}(\Omega)$ and in consequence
$\Delta v_{k}=-Vv_{k}+\lambda v_{k}+G_{k}\in L^{q}(\Omega).$
By [13, Theorem 2.4.2.6, page 241], $v_{k}\in W^{2,q}(\Omega)$. Then we proceed as in the proof of Lemma 2.4 to obtain
$\|v_{k}\|_{W^{2,q}(\Omega)}\leq\mathbf{c}\|(-\Delta+\lambda_{0})v_{k}\|_{L^{q}(\Omega)},$
where $\lambda_{0}=\lambda_{0}(n,\Omega)>0$ is a constant, and consequently
$\|v_{k}\|_{W^{2,q}(\Omega)}$ $\leq\mathbf{c}\|-Vv_{k}+(\lambda+\lambda_{0})v_{k}+G_{k}\|_{L^{q}(\Omega)}$
$\leq\mathbf{c}\left(\|V\|_{L^{n/2}(\Omega)}\|v_{k}\|_{L^{2}(\Omega)}+|\lambda+\lambda_{0}|\|v_{k}\|_{L^{q}(\Omega)}+\|G_{k}\|_{L^{q}(\Omega)}\right).$
From (B.4), $(v_{k})$ is bounded in $H_{0}^{1}(\Omega)$. Hence the last inequalities show that $(v_{k})$ is bounded in $W^{2,q}(\Omega)$ with
(B.5) $\|v_{k}\|_{W^{2,q}(\Omega)}\leq\varkappa\|G_{k}\|_{L^{q}(\Omega)}.$
Here and henceforth, $\varkappa=\varkappa(n,\Omega,V,\lambda)>0$ is a generic constant.
Now as $W^{2,q}(\Omega)$ is reflexive, subtracting if necessary a subsequence, we may assume that $(v_{k})$ converges weakly in $W^{2,p}(\Omega)\cap W_{0}^{1,q}(\Omega)$ to $v\in W^{2,q}(\Omega)\cap W_{0}^{1,q}(\Omega)$. Then $\Delta v+Vv-\lambda v=G$ in the distributional sense and, by using that a norm is weakly lower semi-continuous, we obtain from (B.5)
(B.6) $\|v\|_{W^{2,q}(\Omega)}\leq\varkappa\|G\|_{L^{q}(\Omega)}.$
The function $u=F+v\in W^{2,p}(\Omega)$ is clearly a solution of the BVP (B.1) and inequality (B.2) is a straightforward consequence of inequalities (B.3) and (B.6).
The uniqueness of solutions of the BVP (B.1) follows from the fact that $\lambda$ is not an eigenvalue of $A_{V}$. ∎

### v3

@@
 $\tilde{\mathscr{D}}(V_{1},V_{2})=\sum_{k\geq 1}k^{-2/n}\left[|\lambda_{V_{1}}^{k}-\lambda_{V_{2}}^{k}|+\|\psi_{V_{2}}^{k}-\psi_{V_{1}}^{k}\|_{L^{\tilde{q}}(\Gamma)}\right],$
-$\tilde{\mathscr{D}}_{+}(V_{1},V_{2})=\sum_{k\geq 1}k^{-2/n}\left[|\lambda_{V_{1}}^{k}-\lambda_{V_{2}}^{k}|+\|\phi_{V_{2}}^{k}-\phi_{V_{1}}^{k}\|_{W^{2,q}(\Omega)}\right]$
+$\tilde{\mathscr{D}}_{+}(V_{1},V_{2})=\sum_{k\geq 1}k^{-2/n}\left[|\lambda_{V_{1}}^{k}-\lambda_{V_{2}}^{k}|+\|\phi_{V_{2}}^{k}-\phi_{V_{1}}^{k}\|_{W^{2,\tilde{q}}(\Omega)}\right],$
 $\tilde{\mathscr{W}}=\{(V_{1},V_{2})\in\tilde{\mathscr{V}}\times\tilde{\mathscr{V}};\;|V_{1}-V_{2}|\leq W_{0}\},$
@@
 Let $V\in L^{n/2}(\Omega,\mathbb{R})$ and $\lambda\in\rho(A_{V})$.
-It follows from Theorem B.1 in Appendix B that, for all $\varphi\in W^{2-1/q,q}(\Gamma)$, the BVP (2.7) admits a unique solution $u_{V}(\lambda)(\varphi)\in W^{2,q}(\Omega)$.
+It follows from Theorem B.1 in Appendix B that, for all $\varphi\in W^{2-1/p,p}(\Gamma)$, the BVP (2.7) admits a unique solution $u_{V}(\lambda)(\varphi)\in W^{2,p}(\Omega)$.
 Furthermore, we have
-(2.8) $\|u_{V}(\lambda)(\varphi)\|_{W^{2,q}(\Omega)}\leq\varkappa\|\varphi\|_{W^{2-1/q,q}(\Gamma)},$
+(2.8) $\|u_{V}(\lambda)(\varphi)\|_{W^{2,p}(\Omega)}\leq\varkappa\|\varphi\|_{W^{2-1/p,p}(\Gamma)},$
 where $\varkappa=\varkappa(n,\Omega,V,\lambda)>0$ is a constant.
 Define the family of Dirichlet to Neumann maps $(\Lambda_{V}(\lambda))_{\lambda\in\rho(A_{V})}$ associated with $V\in L^{n/2}(\Omega,\mathbb{R})$ as follows
-$\Lambda_{V}(\lambda):\varphi\in W^{2-1/q,q}(\Gamma)\mapsto\gamma_{1}u_{V}(\lambda)(\varphi)\in W^{1-1/q,q}(\Gamma).$
-By the continuity of the trace map $w\in W^{2,q}(\Omega)\mapsto\gamma_{1}w\in W^{1-1/q,q}(\Gamma)$, it follows from (2.8) that the mapping $\Lambda_{V}(\lambda)$ defines a bounded operator.
+$\Lambda_{V}(\lambda):\varphi\in W^{2-1/p,p}(\Gamma)\mapsto\gamma_{1}u_{V}(\lambda)(\varphi)\in W^{1-1/p,p}(\Gamma).$
+By the continuity of the trace map $w\in W^{2,p}(\Omega)\mapsto\gamma_{1}w\in W^{1-1/p,p}(\Gamma)$, it follows from (2.8) that the mapping $\Lambda_{V}(\lambda)$ defines a bounded operator.
 2.3. Analyticity of solutions
@@
 (2.10) $\|V\phi_{V}^{k}\|_{L^{p}(\Omega)}\leq\mathbf{c}(1+|\lambda_{V}^{k}|)^{1/2}.$
-By using that $\Delta\phi_{V}^{k}=(\lambda_{V}^{k}-V)\phi_{V}^{k}\in L^{q}(\Omega)$, we get $\phi_{k}\in W^{2,q}(\Omega)$.
+By using that $\Delta\phi_{V}^{k}=(\lambda_{V}^{k}-V)\phi_{V}^{k}\in L^{p}(\Omega)\subset L^{q}(\Omega)$ (with continuous embedding), we get $\phi_{k}\in W^{2,q}(\Omega)$.
 In light of $W^{2,q}$ a priori estimate (e.g.
@@
 (2.11) $\|\phi_{k}\|_{W^{2,q}(\Omega)}\leq\mathbf{c}(1+|\lambda_{V}^{k}|).$
-Let $\lambda\in\rho(A_{V})$, $\varphi\in W^{2-1/q,q}(\Gamma)$ and $u:=u_{V}(\lambda)(\varphi)$. Then
-$0=\int_{\Omega}(-\Delta u+(V-\lambda)u)\overline{\phi_{V}^{k}}dx.$
+Let $\lambda\in\rho(A_{V})$, $\varphi\in W^{2-1/p,p}(\Gamma)$ and $u:=u_{V}(\lambda)(\varphi)$. Then
+$0=\int_{\Omega}\left(\nabla u\cdot\nabla\overline{\phi_{V}^{k}}+(V-\lambda)u\overline{\phi_{V}^{k}}\right)dx,$
+where we used that $\nabla u\in L^{2}(\Omega,\mathbb{C}^{n})$ and $u\in L^{p^{\prime}}(\Omega)$.
 Let $(\cdot,\cdot)$ denotes the inner product of $L^{2}(\Omega)$. By performing an integration by parts, we derive from the preceding identity
@@
 is holomorphic and
-(2.14) $u_{V}^{(j)}(\lambda)(\varphi)=-j!\sum_{k\geq 1}\frac{\langle\varphi,\psi_{V}^{k}\rangle}{(\lambda_{V}^{k}-\lambda)^{j+1}}\phi_{V}^{k},\quad j\geq 0.$
-The following notation will be used hereinafter.
-$w^{(j)}:=\frac{d^{j}}{d\lambda^{j}}w.$
-Let $\mathscr{S}$ be the bounded affine function $f\in L^{q}(\Omega)\mapsto\mathscr{S}f:=w\in W^{2,q}(\Omega)$, where $w$ is the unique solution of the BVP
-$\Delta w=f\;\mathrm{in}\;\Omega,\quad u=\varphi\;\mathrm{on}\;\Gamma.$
-By observing that $u_{V}(\lambda)(\varphi)=\mathscr{S}((\lambda-V)u_{V}(\lambda)(\varphi))$, we conclude that $\lambda\in\rho(A_{V})\mapsto u_{V}(\lambda)(\varphi)\in W^{2,q}(\Omega)$ is holomorphic and by induction in $j\geq 1$, we have, where $u(\lambda):=u_{V}(\lambda)(\varphi)$,
-(2.15) $(-\Delta+V-\lambda)u^{(j)}(\lambda)=ju^{(j-1)}(\lambda)\;\mathrm{in}\;\Omega,\quad u^{(j)}(\lambda)=0\;\mathrm{on}\;\Gamma.$
+(2.14) $u_{V}^{(j)}(\lambda)(\varphi)=-j!\sum_{k\geq 1}\frac{\langle\varphi,\psi_{V}^{k}\rangle}{(\lambda_{V}^{k}-\lambda)^{j+1}}\phi_{V}^{k},\quad j\geq 0,$
+where we used the notation
+$w^{(j)}:=\frac{d^{j}}{d\lambda^{j}}w,$
+which we will keep from now on.
+In fact, we have $\lambda\in\rho(A_{V})\mapsto u_{V}(\lambda)(\varphi)\in W^{2,p}(\Omega)$ is holomorphic as shown by the following proposition.
+Proposition 2.2.
+For all $V\in\mathscr{V}$, $\varphi\in W^{2-1/p,p}(\Gamma)$, $\lambda\in\rho(A_{V})\mapsto u_{V}(\lambda)(\varphi)\in W^{2,p}(\Omega)$ is holomorphic and $u^{(1)}(\lambda):=u_{V}^{(1)}(\lambda)(\varphi)$ is the solution of the BVP
+(2.15) $(-\Delta+V-\lambda)u^{(1)}(\lambda)=u_{V}(\lambda)(\varphi)\;\mathrm{in}\;\Omega,\quad u=0\;\mathrm{on}\;\Gamma,$
+Proof.
+Let $V\in\mathscr{V}$, $\lambda\in\rho(A_{V})$ and $\mu\in B(0,\mathrm{dist}(\lambda,\sigma(A_{V})/2))$.
+From Theorem B.1 and its proof, the constant $\mathbf{c}_{\lambda+\mu}$ in (B.5) can be chosen such that
+$\mathbf{c}_{\lambda}:=\sup_{\mu\in B(0,\mathrm{dist}(\lambda,\sigma(A_{V})/2)}\mathbf{c}_{\lambda+\mu}<\infty.$
+Since $v=u(\lambda+\mu)(\varphi)-u(\lambda)(\varphi)$ is the solution of the BVP
+$(-\Delta+V-(\lambda+\mu))v=\mu u(\lambda)(\varphi)\;\mathrm{in}\;\Omega,\quad v=0\;\mathrm{on}\;\Gamma,$
+we obtain from (B.5)
+(2.16) $\|u(\lambda+\mu)(\varphi)-u(\lambda)(\varphi)\|_{W^{2,p}(\Omega)}\leq\mathbf{c}_{\lambda}|\mu|\|u(\lambda)(\varphi)\|_{L^{p}(\Omega)}.$
+Hence, $z\in\rho(A_{V})\mapsto u_{V}(z)(\varphi)\in W^{2,p}(\Omega)$ is continuous.
+Next, let $u^{(1)}(\lambda)$ be the solution of (2.15) and set
+$w:=u(\lambda+\mu)(\varphi)-u(\lambda)(\varphi)-\mu u^{(1)}(\lambda).$
+We verify that $w$ is the solution of the BVP
+$(-\Delta+V-\lambda)w=\mu[u(\lambda+\mu)(\varphi)-u(\lambda)(\varphi)]\;\mathrm{in}\;\Omega,\quad w=0\;\mathrm{on}\;\Gamma,$
+Applying (B.5) and then (2.16), we obtain
+$\|u(\lambda+\mu)(\varphi)-u(\lambda)(\varphi)-\mu u^{(1)}(\lambda)\|_{W^{2,p}(\Omega)}\leq\mathbf{c}_{\lambda}^{2}|\mu|^{2}\|u(\lambda)(\varphi)\|_{L^{p}(\Omega)}.$
+This inequality leads to the expect result. ∎
+Since $W^{2,p}(\Omega)$ is continuously embedded in $W^{2,p}(\Omega)$, the map $\lambda\in\rho(A_{V})\mapsto u_{V}(\lambda)(\varphi)\in W^{2,q}(\Omega)$ is also holomorphic for all $\varphi\in W^{2-1/p,p}(\Gamma)$.
+This observation will be used in the following.
+Let $V\in\mathscr{V}$, $\varphi\in W^{2-1/p,p}(\Gamma)$ and $\lambda\in\rho(A_{V})$. By induction in $j\geq 1$, we derive from (2.15)
+(2.17) $(-\Delta+V-\lambda)u_{V}^{(j)}(\lambda)(\varphi)=ju_{V}^{(j-1)}(\lambda)(\varphi)\;\mathrm{in}\;\Omega,\quad u^{(j)}(\lambda)=0\;\mathrm{on}\;\Gamma.$
 2.4. Resolvent estimates
@@
 Let
-$\mathcal{B}=W^{2-1/q,q}(\Gamma)\cap L^{q^{\prime}}(\Gamma)$
+$\mathcal{B}=W^{2-1/p,p}(\Gamma)\cap L^{q^{\prime}}(\Gamma)$
 be endowed with the norm
-$\|\varphi\|:=\|\varphi\|_{W^{2-1/q,q}(\Gamma)}+\|\varphi\|_{L^{q^{\prime}}(\Gamma)},\quad\varphi\in\mathcal{B}.$
+$\|\varphi\|:=\|\varphi\|_{W^{2-1/p,p}(\Gamma)}+\|\varphi\|_{L^{q^{\prime}}(\Gamma)},\quad\varphi\in\mathcal{B}.$
 As in the previous subsection, $\mathbf{c}=\mathbf{c}(n,\Omega,V_{0})>0$ still denotes a generic constant.
@@
 $\mathscr{D}(V_{1},V_{2}):=\sum_{k\geq 1}k^{-2/n}\left[|\lambda_{V_{1}}^{k}-\lambda_{V_{2}}^{k}|+\|\psi_{V_{2}}^{k}-\psi_{V_{1}}^{k}\|_{L^{q}(\Gamma)}\right]<\infty,$
-Recall also that $\mathcal{B}:=W^{2-1/q,q}(\Gamma)\cap L^{q^{\prime}}(\Gamma)$ is equipped with the norm
-$\|\varphi\|=\|\varphi\|_{W^{2-1/q,q}(\Gamma)}+\|\varphi\|_{L^{q^{\prime}}(\Gamma)},\quad\varphi\in\mathcal{B}.$
+Recall also that $\mathcal{B}:=W^{2-1/p,p}(\Gamma)\cap L^{q^{\prime}}(\Gamma)$ is equipped with the norm
+$\|\varphi\|=\|\varphi\|_{W^{2-1/p,p}(\Gamma)}+\|\varphi\|_{L^{q^{\prime}}(\Gamma)},\quad\varphi\in\mathcal{B}.$
 In the remaining part of this section, $\mathbf{c}=\mathbf{c}(n,\Omega,V_{0},W_{0},t)>0$ will denote a generic constant.
@@
 (3.3) $\mathbf{c}\|\Lambda_{V_{1}}(\lambda_{\tau})-\Lambda_{V_{2}}(\lambda_{\tau})\|_{L^{q}(\Gamma)}\leq\tau^{-3}\|\varphi\|$
-$\hskip 85.35826pt+\tau^{\sigma}\int_{0}^{1}\|\Lambda_{V_{1}}(\mu_{\tau}^{s}(\varphi))-\Lambda_{V_{2}}(\mu_{\tau}^{s}(\varphi))\|_{L^{q}(\Gamma)}ds.$
+$\hskip 85.35826pt+\tau^{\sigma}\int_{0}^{1}\|\Lambda_{V_{1}}^{(1)}(\mu_{\tau}^{s}(\varphi))-\Lambda_{V_{2}}^{(1)}(\mu_{\tau}^{s}(\varphi))\|_{L^{q}(\Gamma)}ds.$
 In light of (2.5), we find an integer $\ell_{\ast}=\ell_{\ast}(n,\Omega,V_{0})>0$ and a constant $\mathbf{c}_{0}=(n,\Omega,V_{0})>0$ such that
@@
 In light of (3.7), a combination of (3.11) and (3.12) gives
-(3.13) $\|\gamma_{1}v_{V_{1}}^{(1)}(\mu_{\tau}^{s})(\varphi)-\gamma_{1}v_{V_{2}}^{(1)}(\mu_{\tau}^{s})(\varphi)\|_{L^{q}(\Gamma)}\leq\mathbf{c}\mathscr{D}(V_{1},V_{2})\|\varphi\|,$
+(3.13) $\|\gamma_{1}v_{V_{1}}^{(1)}(\mu_{\tau}^{s})(\varphi)-\gamma_{1}v_{V_{2}}^{(1)}(\mu_{\tau}^{s})(\varphi)\|_{L^{q}(\Gamma)}\leq\mathbf{c}\mathscr{D}(V_{1},V_{2})\|\varphi\|.$
 We proceed similarly to prove
@@
 In light of the definition of $k_{\tau}$, we verify that
-(3.14) $\|\gamma_{1}w_{V_{1}}^{(1)}(\mu_{\tau}^{s})(\varphi)-\gamma_{1}w_{V_{2}}^{(1)}(\mu_{\tau}^{s})(\varphi)\|_{L^{q}(\Gamma)}\leq\mathbf{c}\tau^{n+2}\mathscr{D}(V_{1},V_{2})\|\varphi\|$
+(3.14) $\|\gamma_{1}w_{V_{1}}^{(1)}(\mu_{\tau}^{s})(\varphi)-\gamma_{1}w_{V_{2}}^{(1)}(\mu_{\tau}^{s})(\varphi)\|_{L^{q}(\Gamma)}\leq\mathbf{c}\tau^{n+2}\mathscr{D}(V_{1},V_{2})\|\varphi\|.$
 Using (3.13) and (3.14) in (3.3), we obtain
@@
 Appendix B BVPs with unbounded lower order term
+We will use the following lemma, the proof of which is quite similar to that of Lemma 2.2.
+Lemma B.1.
+Let $V\in\mathscr{V}$, $\lambda\in\rho(A_{V})$. Then
+(B.1) $\|R_{V}(\lambda)f\|_{L^{2}(\Omega)}\leq\frac{1}{\mathrm{dist}(\lambda,\sigma(A_{V}))}\|f\|_{L^{2}(\Omega)},\quad f\in L^{2}(\Omega),$
+(B.2) $\|R_{V}(\lambda)f\|_{H_{0}^{1}(\Omega)}\leq\mathbf{c}_{\lambda}\|f\|_{L^{2}(\Omega)},\quad f\in L^{2}(\Omega),$
+(B.3) $\|R_{V}(\lambda)f\|_{H_{0}^{1}(\Omega)}\leq\mathbf{c}_{\lambda}\|f\|_{L^{p}(\Omega)},\quad\quad f\in L^{p}(\Omega).$
+Here, $\mathbf{c}_{\lambda}=\mathbf{c}_{\lambda}(n,\Omega,V,\lambda)>0$ is a constant.
 We consider the BVP
@@
 Let $V\in L^{n/2}(\Omega,\mathbb{R})$ and $\lambda\in\rho(A_{V})$.
-For all $f\in L^{q}(\Omega)$ and $\varphi\in W^{2-1/q,q}(\Gamma)$ the BVP (B.1) has a unique solution $u=u_{V}(\lambda)(\varphi,f)\in W^{2,q}(\Omega)$.
-In addition, there exists a constant $\varkappa=\varkappa(n,\Omega,V,\lambda)>0$ such that
-(B.2) $\|u_{V}(\lambda)(\varphi,f)\|_{W^{2,q}(\Omega)}\leq\varkappa\left(\|f\|_{L^{q}(\Omega)}+\|\varphi\|_{W^{2-1/q,q}(\Gamma)}\right).$
+For all $f\in L^{p}(\Omega)$ and $\varphi\in W^{2-1/p,p}(\Gamma)$ the BVP (B.4) has a unique solution $u=u_{V}(\lambda)(\varphi,f)\in W^{2,p}(\Omega)$.
+In addition, there exists a constant $\mathbf{c}_{\lambda}=\mathbf{c}_{\lambda}(n,\Omega,V,\lambda)>0$ such that
+(B.5) $\|u_{V}(\lambda)(\varphi,f)\|_{W^{2,p}(\Omega)}\leq\mathbf{c}_{\lambda}\left(\|f\|_{L^{p}(\Omega)}+\|\varphi\|_{W^{2-1/p,p}(\Gamma)}\right).$
 Proof.
-In this proof, $\mathbf{c}_{0}=\mathbf{c}_{0}(n,\Omega)>0$ and $\mathbf{c}=\mathbf{c}(n,\Omega,V_{0})>0$ are generic constants.
-Let $F\in W^{2,q}(\Omega)$ so that $\gamma_{0}F=\varphi$ and
-$\|F\|_{W^{2,q}(\Omega)}\leq\mathbf{c}_{0}\|\varphi\|_{W^{2-1/q,q}(\Gamma)}.$
-If $G=-(-\Delta+V-\lambda)F+f$ then $G\in L^{q}(\Omega)$ and
-(B.3) $\|G\|_{L^{q}(\Omega)}\leq\mathbf{c}\left(\|f\|_{L^{q}(\Omega)}+\|\varphi\|_{W^{2-1/q,q}(\Gamma)}\right).$
-By density, we can find a sequence $(G_{k})$ in $L^{2}(\Omega)$ converging in $L^{q}(\Omega)$ to $G$. Set for any $k$
-(B.4) $v_{k}=R_{V}(\lambda)G_{k}.$
-Then $v_{k}\in H_{0}^{1}(\Omega)$ and in consequence
-$\Delta v_{k}=-Vv_{k}+\lambda v_{k}+G_{k}\in L^{q}(\Omega).$
-By [13, Theorem 2.4.2.6, page 241], $v_{k}\in W^{2,q}(\Omega)$. Then we proceed as in the proof of Lemma 2.4 to obtain
-$\|v_{k}\|_{W^{2,q}(\Omega)}\leq\mathbf{c}\|(-\Delta+\lambda_{0})v_{k}\|_{L^{q}(\Omega)},$
+In this proof, $\mathbf{c}_{0}=\mathbf{c}_{0}(n,\Omega)>0$, $\mathbf{c}=\mathbf{c}(n,\Omega,V)>0$ and $\mathbf{c}_{\lambda}=\mathbf{c}_{\lambda}(n,\Omega,V,\lambda)$ are generic constants.
+Let $F\in W^{2,p}(\Omega)$ so that $\gamma_{0}F=\varphi$ and
+$\|F\|_{W^{2,p}(\Omega)}\leq\mathbf{c}_{0}\|\varphi\|_{W^{2-1/p,p}(\Gamma)}.$
+If $G=-(-\Delta+V-\lambda)F+f$ then $G\in L^{p}(\Omega)$ and
+(B.6) $\|G\|_{L^{p}(\Omega)}\leq\mathbf{c}_{\lambda}\left(\|f\|_{L^{p}(\Omega)}+\|\varphi\|_{W^{2-1/p,p}(\Gamma)}\right).$
+It follows from (B.3) that $v=R_{V}(\lambda)f\in H_{0}^{1}(\Omega)$ and
+(B.7) $\|v\|_{H_{0}^{1}(\Omega)}\leq\mathbf{c}_{\lambda}\|G\|_{L^{p}(\Omega)}.$
+Furthermore, we have
+$\Delta v=-Vv+\lambda v+G\in L^{p}(\Omega).$
+By [13, Theorem 2.4.2.6, page 241], we obtain $v\in W^{2,p}(\Omega)$. Then we proceed as in the proof of Lemma 2.4 to obtain
+$\|v\|_{W^{2,p}(\Omega)}\leq\mathbf{c}\|(-\Delta+\lambda_{0})v\|_{L^{p}(\Omega)},$
 where $\lambda_{0}=\lambda_{0}(n,\Omega)>0$ is a constant, and consequently
-$\|v_{k}\|_{W^{2,q}(\Omega)}$ $\leq\mathbf{c}\|-Vv_{k}+(\lambda+\lambda_{0})v_{k}+G_{k}\|_{L^{q}(\Omega)}$
-$\leq\mathbf{c}\left(\|V\|_{L^{n/2}(\Omega)}\|v_{k}\|_{L^{2}(\Omega)}+|\lambda+\lambda_{0}|\|v_{k}\|_{L^{q}(\Omega)}+\|G_{k}\|_{L^{q}(\Omega)}\right).$
-From (B.4), $(v_{k})$ is bounded in $H_{0}^{1}(\Omega)$. Hence the last inequalities show that $(v_{k})$ is bounded in $W^{2,q}(\Omega)$ with
-(B.5) $\|v_{k}\|_{W^{2,q}(\Omega)}\leq\varkappa\|G_{k}\|_{L^{q}(\Omega)}.$
-Here and henceforth, $\varkappa=\varkappa(n,\Omega,V,\lambda)>0$ is a generic constant.
-Now as $W^{2,q}(\Omega)$ is reflexive, subtracting if necessary a subsequence, we may assume that $(v_{k})$ converges weakly in $W^{2,p}(\Omega)\cap W_{0}^{1,q}(\Omega)$ to $v\in W^{2,q}(\Omega)\cap W_{0}^{1,q}(\Omega)$.
-Then $\Delta v+Vv-\lambda v=G$ in the distributional sense and, by using that a norm is weakly lower semi-continuous, we obtain from (B.5)
-(B.6) $\|v\|_{W^{2,q}(\Omega)}\leq\varkappa\|G\|_{L^{q}(\Omega)}.$
-The function $u=F+v\in W^{2,p}(\Omega)$ is clearly a solution of the BVP (B.1) and inequality (B.2) is a straightforward consequence of inequalities (B.3) and (B.6).
-The uniqueness of solutions of the BVP (B.1) follows from the fact that $\lambda$ is not an eigenvalue of $A_{V}$. ∎
+$\|v\|_{W^{2,q}(\Omega)}$ $\leq\mathbf{c}_{0}\|-Vv+(\lambda+\lambda_{0})v+G\|_{L^{p}(\Omega)}$
+$\leq\mathbf{c}_{0}\left(\|V_{0}\|_{L^{n/2}(\Omega)}\|v\|_{L^{p^{\prime}}(\Omega)}+|\lambda+\lambda_{0}|\|v\|_{L^{p}(\Omega)}+\|G\|_{L^{p}(\Omega)}\right).$
+The last inequality, combined with (B.7), gives
+$\|v\|_{W^{2,p}(\Omega)}\leq\mathbf{c}_{\lambda}\|G\|_{L^{p}(\Omega)}.$
+That is, we have
+(B.8) $\|v\|_{W^{2,p}(\Omega)}\leq\mathbf{c}_{\lambda}(\|f\|_{L^{p}(\Omega)}+\|\varphi\|_{W^{2-1/p,p}(\Gamma)}).$
+We verify that $u=F+v\in W^{2,p}(\Omega)$ is a solution of the BVP (B.4) and inequality (B.5) follows from (B.8). This solution is unique because $\lambda$ is not an eigenvalue of $A_{V}$. ∎